### v5

# arXiv:2406.18500v5[math.AP] 01 Mar 2026

Abstract
In this paper, we study linear backward parabolic SPDEs in bounded domains and present new a priori estimates for their weak solutions. Inspired by the seminal work of Y. Hu, J. Ma and J. Yong from 2002 on strong solutions, we establish $L^{p}$-estimates requiring minimal assumptions on the regularity of the coefficients, the terminal data, and the external force. Our approach relies on direct, constructive, and quantitative arguments, adapted from known methods in the theory of SPDEs to this setting. In particular, we develop a new Itô’s formula for the $L^{p}$-norm of the backward solution, tailored to this setting and extending the classical result in the $L^{2}$-framework. This formula is then used to improve further the regularity of the first component of the solution up to $L^{\infty}$. We also present two applications: a local existence result for a semilinear equation without imposing any growth condition on the nonlinear term, and a novel local controllability result for semilinear backward SPDEs that partially resolves an open problem in the field.
Keywords: Backward SPDEs, weak solutions, Itô’s formula, controllability, Banach fixed-point.
1 Introduction
1.1 Notation and main results
Backward stochastic partial differential equations (BSPDEs, for short) arise in many applications of probability theory and stochastic processes. In the literature, BSPDEs can be found in various contexts, including control under incomplete information (see [BEN83, TAN98]), as adjoint equations in optimal control (see e.g. [NN90, HP91, ZHO93]), in controllability problems (see e.g. [BRT03, TZ09, HLP23]), in mathematical finance for the formulation of the stochastic Feynman-Kac formula and the Black-Scholes equation (see [MY97, MY99]), among other applications.
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a complete probability space on which is defined a one-dimensional standard Brownian motion $W=\{W(t):t\geq 0\}$ and let $\{\mathcal{F}_{t}\}_{t\geq 0}$ be the natural filtration generated by $W$, augmented by all the $\mathbb{P}$-null sets in $\mathcal{F}$. For any given (deterministic) time $T>0$, let us consider the following linear BSPDE
$\begin{cases}\textnormal{d}{y}=-\left(\Delta y+\alpha y+\beta Y+F\right)\textnormal{d}{t}+Y\textnormal{d}{W}(t)&\text{in }(0,T)\times\mathcal{O},\\ y=0&\text{on }(0,T)\times\partial\mathcal{O},\\ y(T)=y_{T}&\text{in }\mathcal{O},\end{cases}$ (1)
where $\mathcal{O}\subset\mathbb{R}^{d}$ ($d\geq 1$) is an open and bounded set with a $C^{2}$ boundary $\partial\mathcal{O}$. In (1), the coefficients $\alpha$ and $\beta$, the forcing term $F$ and the terminal datum $y_{T}$ are suitable random fields verifying appropriate measurability and regularity conditions.
A solution to (1) consists of a pair of random fields $(y,Y)$ that are $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted and satisfy (1) in a specified sense. At this point, it is important to note a fundamental difference from standard SPDEs: the noise term in a BSPDE like (1) arises from the intrinsic randomness in the coefficients and terminal data—making it endogenous—and comes from martingale representation theorems.
A classical question for BSPDEs consists in understanding the notion of the solution (for instance strong, mild, or weak in the PDE sense) and determining the best regularity available for the given data of the problem. This issue has been studied in the literature in different contexts and frameworks. In [HP91, Lv01], using semigroup theory, the authors establish the existence and uniqueness of mild solutions for abstract backward evolution equations in the $L^{2}$ and $L^{p}$ settings, respectively. In [HMY02, DOK12, DT12], the problem of existence and regularity of strong solutions (i.e., solutions that can be evaluated point-wisely in the spatial variable) is studied in bounded and/or unbounded domains. In the context of weak solutions (i.e., the PDE satisfies a scalar product equation), the works [DQT12, DTZ13] tackle the existence and regularity of solutions posed in the whole space within the $L^{p}$-setting and in higher order Sobolev spaces $W^{m,p}$ ($m\geq 1$, $p\geq 2$), respectively. Finally, existence and regularity of solutions in Hölder spaces are established in [TW16].
In this direction, the first goal of this paper is to establish $L^{p}$- and $L^{\infty}$-estimates for weak solutions to (1) (see ˜1.1 below) without imposing additional regularity and/or differentiability assumptions on the coefficients, the terminal data or the source term. Our aim is to use natural hypotheses corresponding to the desired function space: for instance, obtaining $L^{p}$ estimates requires $L^{p}$ terminal datum and source term, and similarly, $L^{\infty}$ solutions are obtained with $L^{\infty}$ data.
To make this precise, we begin by introducing the following notations. In what follows, we denote $\{\mathcal{F}_{t}\}_{t\geq 0}$ by $\mathbb{F}$ unless we want to emphasize a precise $\mathcal{F}_{t}$. Let $(X,\|\cdot\|_{X})$ be a Banach space, for any $p\in[2,+\infty)$ and any $s\in\{0,T\}$, we denote by $L^{p}_{\mathcal{F}_{s}}(\Omega;X)$ the set of all $\mathcal{F}_{s}$-measurable $X$-valued random variables $\xi:\Omega\to X$ such that ${\mathbb{E}}(|\xi|_{X}^{p})<+\infty$. For any $p,q\in[2,\infty)$ we define the spaces
[…]
endowed with their natural norms. In a clear and analogous way, we define the spaces (2) and (3) when $p$ and/or $q$ equal to $\infty$. When $p=q$ with $p\in[2,+\infty)$, from Fubini’s theorem we have that $L_{\mathbb{F}}^{p}(\Omega;L^{p}(0,T;X))=L_{\mathbb{F}}^{p}(0,T;L^{p}(\Omega;X))$ and the norms coincide and for simplicity we just write $L_{\mathbb{F}}^{p}(0,T;X)$. When $p=q=\infty$, we also have that $L_{\mathbb{F}}^{\infty}(\Omega;L^{\infty}(0,T;X))=L_{\mathbb{F}}^{\infty}(0,T;L^{\infty}(\Omega;X))$ and the norms coincide. Lastly, by $L_{\mathbb{F}}^{p}(\Omega;C([0,T];X))$ we denote the Banach space consisting of all $X$-valued $\mathbb{F}$-adapted continuous processes $\psi(\cdot)$ such that $\mathbb{E}\left(\left\lVert\psi\right\rVert_{C([0,T];X)}^{p}\right)<+\infty$, also equipped with the canonical norm.
In the remainder of this document, we will make the following instrumental assumption on the coefficients of system (1)
  1. (H)
    $\alpha,\beta\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))$.
We recall the following notion of solution for (1).
Definition 1.1.
A pair of random fields $(y,Y)$ is called a weak solution to (1) if
  1. 1)
    $(y,Y)$ is $L^{2}(\mathcal{O})\times L^{2}(\mathcal{O})$-valued and $\mathcal{F}_{t}$-measurable for each $t\in[0,T]$,
  2. 2)
    $(y,Y)\in\left[L^{2}_{\mathbb{F}}(\Omega;C([0,T]);L^{2}(\mathcal{O}))\bigcap L^{2}_{\mathbb{F}}(0,T;H_{0}^{1}(\mathcal{O}))\right]\times L^{2}_{\mathbb{F}}(0,T;L^{2}(\mathcal{O}))$, and
  3. 3)
    for any $t\in[0,T]$ and $\phi\in H_{0}^{1}(\mathcal{O})$ it holds
    $(y(t),\phi)_{L^{2}(\mathcal{O})}$ $=(y_{T},\phi)_{L^{2}(\mathcal{O})}-\int_{t}^{T}(\nabla y,\nabla\phi)_{L^{2}(\mathcal{O})}\textnormal{d}{s}$
    $\quad+\int_{t}^{T}(\alpha y+\beta Y+F,\phi)_{L^{2}(\mathcal{O})}\textnormal{d}{s}-\int_{t}^{T}(Y,\phi)_{L^{2}(\mathcal{O})}\textnormal{d}{W}(s),\quad\textnormal{a.s.}$
The following result ensures the existence and uniqueness of such weak solutions.
Theorem 1.2.
Let $\alpha,\beta$ be given coefficients satisfying (H). If $y_{T}\in L^{2}_{\mathcal{F}_{T}}(\Omega;L^{2}(\mathcal{O}))$ and $F\in L^{2}_{\mathbb{F}}(0,T;L^{2}(\mathcal{O}))$, then (1) admits a unique weak solution in the sense of ˜$1.$ Moreover, the following energy estimate holds
${\mathbb{E}}\left(\sup_{t\in[0,T]}\|y(t)\|^{2}_{L^{2}(\mathcal{O})}\right)+{\mathbb{E}}\left(\int_{0}^{T}\int_{\mathcal{O}}|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}t\right)+{\mathbb{E}}\left(\int_{0}^{T}\int_{\mathcal{O}}|Y|^{2}\,\textnormal{d}x\textnormal{d}t\right)$
$\quad\leq C{\mathbb{E}}\left(\int_{\mathcal{O}}|y_{T}|^{2}\,\textnormal{d}x+\int_{0}^{T}\int_{\mathcal{O}}|F|^{2}\,\textnormal{d}x\textnormal{d}t\right),$ (4)
for a positive constant $C>0$ independent of $y_{T}$ and $F$.
This theorem is well-known in the literature and the proof can be derived in two different ways: by a duality analysis similar to [ZHO92] or by a more direct and constructive approach relying on Galerkin method (see, e.g., [GAO18, Proposition 2.1] or [SY09, Section 3]).
Our first main result says that the integrability of the weak solution $(y,Y)$ can be improved by taking more regular data. More precisely, we have the following.
Theorem 1.3.
Let $(y,Y)$ be a weak solution to (1). Assume that $y_{T}\in L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))$ and $F\in L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))$ for some $p\in[2,+\infty)$. Then,
$(y,Y)\in L^{p}_{\mathbb{F}}(\Omega;C([0,T];L^{p}(\mathcal{O})))\times L^{p}_{\mathbb{F}}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O})))$
and satisfies
${\mathbb{E}}\left(\sup_{t\in[0,T]}\|y(t)\|^{p}_{L^{p}(\mathcal{O})}\right)+{\mathbb{E}}\left[\left(\int_{0}^{T}\|Y(t)\|_{L^{2}(\mathcal{O})}^{2}\textnormal{d}t\right)^{p/2}\right]$
$\qquad\leq Ce^{CT}{\mathbb{E}}\left(\int_{\mathcal{O}}|y_{T}|^{p}\,\textnormal{d}x+\int_{0}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}t\right),$ (5)
for a constant $C>0$ only depending on $\mathcal{O}$, $\alpha$, $\beta$ and $p$.
Our second result says that we can go up to $L^{\infty}$ for the first component $y$ of the solution $(y,Y)$ to (1). In more detail, we have the following.
Theorem 1.4.
Let $(y,Y)$ be a weak solution to (1). Assume that $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))$ and $F\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))$. Then the process $y$ taken from the solution $(y,Y)$ to (1) belongs to $L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))$ and satisfies
$\operatorname*{ess\,sup}_{(\omega,t)\in\Omega\times[0,T]}\|y(t)\|_{L^{\infty}(\mathcal{O})}\leq\exp(CT)\left(\left\lVert y_{T}\right\rVert_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}+\left\lVert F\right\rVert_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\right),$ (6)
for some $C>0$ only depending on $\mathcal{O}$, $\alpha$ and $\beta$.
Theorems 1.3 and 1.4 should be compared with [HMY02, Theorem 3.2], where similar results are presented for one-dimensional BSPDEs. However, our contributions differ in several ways. Firstly, while [HMY02] is restricted to the one-dimensional case, our results are established for BSPDEs in any dimension. Secondly, we address weak solutions rather than strong solutions. This distinction is important since weak solutions are less regular in the spatial variable and cannot be evaluated pointwise, preventing the direct application of Itô’s formula to the function $y^{2p}$ as in [HMY02]. Instead, we establish an Ito’s formula for the $L^{p}$-norm in the backward case (see Proposition˜2.1 below), using some ideas from [DG15] for SPDEs, which involve direct and constructive arguments, like truncations and limit procedures for a suitable approximation of the $L^{p}$-norm (see Sections 2.1 and 2.2). Finally, our work considers coefficients that are merely bounded, as opposed to some differentiability conditions imposed in [HMY02, eq. (2.4)]. Similarly, we relax the conditions on the initial datum and the source term, which are less restrictive compared to those required by [HMY02, eqs. (2.2)–(2.3)].
As in [HMY02], we also obtain a uniform bound on the first component of the solution of the BSPDE (see Theorem 1.4), which initially can be surprising due to the presence of the stochastic integral. Nevertheless, as mentioned before, the noise term in (1) is endogenous to the equation, and as noted in [DT12], this special feature makes the theory for these equations closer to deterministic PDEs than to SPDEs.
Remark 1.5.
We also note that our approach provides an improved dependence on the variable $\omega\in\Omega$ for the process $Y$; see [HMY02, Theorem 3.2] together with Theorem˜1.3. In contrast, within the $L^{\infty}$ framework of Theorem˜1.4 no improvement in the regularity of $Y$ is obtained. We refer to Section˜6 for additional discussion related to the potential optimality of these results.
Remark 1.6.
The proof of Theorem˜1.4 is obtained by refining some of the estimates derived in the proof of Theorem˜1.3. In particular, we establish a version of (5) with constants that are uniform in $p$, and then pass to the limit as $p\to\infty$.
Alternatively, one might consider using comparison principle arguments to derive the estimate in Theorem˜1.4; see, for instance, [DT12, Theorem 5.1]. That result is stated under assumptions that guarantee the existence of strong solutions and require some differentiability of the initial data. Whether these conditions can be relaxed to cover our setting remains unclear. We therefore choose to follow a different, more quantitative route, which stays within the weak solution framework and naturally yields the $L^{p}$-estimates needed in our applications below.
1.2 Applications
Once we have established the regularity of the BSPDE (1), we are interested in studying the following applications:
  1. 1.
    Local existence of semilinear equations.
  2. 2.
    Controllability for linear and semilinear systems.
The first application is concerned about the local existence of solutions to the semilinear system
$\begin{cases}\textnormal{d}{y}=-\left(\Delta y+\alpha y+\beta Y+f(y)\right)\textnormal{d}{t}+Y\textnormal{d}{W}(t)&\text{in }(0,T)\times\mathcal{O},\\ y=0&\text{on }(0,T)\times\partial\mathcal{O},\\ y(T)=y_{T}&\text{in }\mathcal{O},\end{cases}$ (7)
where we assume that
  1. (F)
    $f\in C^{\infty}(\mathbb{R})$ such that $f(0)=0$.
A pair of random fields $(y,Y)$ is called a weak solution to (7) if
  1. 1)
    $(y,Y)$ is $L^{2}(\mathcal{O})\times L^{2}(\mathcal{O})$-valued and $\mathcal{F}_{t}$-measurable for each $t\in[0,T]$,
  2. 2)
    $(y,Y)\in\left[L^{2}_{\mathbb{F}}(\Omega;C([0,T]);L^{2}(\mathcal{O}))\bigcap L^{2}_{\mathbb{F}}(0,T;H_{0}^{1}(\mathcal{O}))\right]\times L^{2}_{\mathbb{F}}(0,T;L^{2}(\mathcal{O}))$, and
  3. 3)
    for any $t\in[0,T]$ and $\phi\in H_{0}^{1}(\mathcal{O})$ it holds
    $(y(t),\phi)_{L^{2}(\mathcal{O})}$ $=(y_{T},\phi)_{L^{2}(\mathcal{O})}-\int_{t}^{T}(\nabla y,\nabla\phi)_{L^{2}(\mathcal{O})}\textnormal{d}{s}$
    $\quad+\int_{t}^{T}(\alpha y+\beta Y+f(y),\phi)_{L^{2}(\mathcal{O})}\textnormal{d}{s}-\int_{t}^{T}(Y,\phi)_{L^{2}(\mathcal{O})}\textnormal{d}{W}(s),\quad\textnormal{a.s.}$
To state our result, let us define the functional space $\mathscr{X}:=L^{2}_{\mathbb{F}}(\Omega;C([0,T;]L^{2}(\mathcal{O})))\cap L^{2}_{\mathbb{F}}(\Omega;L^{2}(0,T;H_{0}^{1}(\mathcal{O})))\cap L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))$ and for any $p\in[2,+\infty)$ we set
$\mathscr{Y}_{p}:=\left\{(y,Y)\in\mathscr{X}\times L^{p}_{\mathbb{F}}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O})))\right\}$ (8)
endowed with its natural norm, that is, $\|(y,Y)\|_{\mathscr{Y}_{p}}=\|y\|_{\mathscr{X}}+\|Y\|_{L^{p}_{\mathbb{F}}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O})))}.$
Theorem 1.7.
Let $T>0$, $p\in[2,+\infty)$ and assume that (F) holds. There exists $\delta>0$ only depending on $\mathcal{O}$, $T$, $p$ and $f$ such that for any $y_{0}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))$ satisfying $\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\leq\delta$, there is a unique weak solution $(y,Y)$ to (7) that belongs to $\mathscr{Y}_{p}$.
The proof of Theorem˜1.7 relies on a standard Banach fixed-point procedure and exploits the boundedness of the solution provided by Theorem˜1.4 to prove the contractivity of a suitable nonlinear map.
Note that we do not impose any growth or sign conditions on the nonlinear term. Thanks to the higher regularity already established for the solution, we can apply a suitable Taylor expansion to the nonlinear term, which allows us to bypass the usual difficulties associated with low regularity (see Section˜4 and Lemma˜A.3). In this way, we obtain a local well-posedness result for (7) for sufficiently small terminal data. This avoids the standard global Lipschitz assumptions commonly used in the literature (see, e.g., [HP91, HMY02] or [LZ02, Chapter 4]) and provides a more flexible framework for handling semilinear systems.
Our second application is concerned about the controllability of (7). Controllability is a qualitative property of dynamical systems that refers to the ability to steer the system from any initial state to any desired final state within a finite time period, using an appropriate control input. To fix ideas, let us consider the linear control system
$\begin{cases}\textnormal{d}{y}=-\left(\Delta y+\alpha y+\beta Y+\chi_{\mathcal{O}_{0}}h\right)\textnormal{d}{t}+Y\textnormal{d}{W}(t)&\text{in }(0,T)\times\mathcal{O},\\ y=0&\text{on }(0,T)\times\partial\mathcal{O},\\ y(T)=y_{T}&\text{in }\mathcal{O},\end{cases}$ (9)
where $h$ is an external control force localized on a set $\mathcal{O}_{0}\subset\mathcal{O}$. In this part we assume that $\mathcal{O}$ is connected.
Due to Theorems 1.2, 1.3 and 1.4, system (9) is well-posed for any terminal datum $y_{T}\in L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))$ and any function $h\in L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}_{0}))$ with $p\in[2,+\infty]$. This motivates us to find controls $h$ such that the following controllability condition is fulfilled.
Definition 1.8.
Let $p\in[2,\infty]$. System (9) is said to be null-controllable in $L^{p}$ if for any terminal datum $y_{T}\in L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))$, there exists a control $h\in L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}_{0}))$ such that the corresponding weak solution $(y,Y)$ of (9) satisfies
$y(0,\cdot)=0\quad\textnormal{a.s.}$
In the case $p=2$, this problem has been studied extensively, most notably in [BRT03] and [TZ09]. Using duality arguments, the controllability for (9) is established via Carleman estimates, which have been adapted from their classical deterministic form (see [FI96]) to the stochastic setting. However, to our knowledge, there are no results in the literature addressing the case when $p>2$. To bridge this gap, we present a result covering the full range $p\in(2,+\infty]$ for a simplified yet interesting case.
Proposition 1.9.
Let $p\in(2,+\infty]$ and assume that $\beta\equiv 0$. Then, system (9) is null-controllable in $L^{p}$. Moreover, if $T\in(0,1)$, the following estimate on the control holds
$\|h\|_{L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))}\leq\exp(C/T)\|y_{T}\|_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))},$ (10)
for some $C>0$ depending at most on $\mathcal{O}$, $\mathcal{O}_{0}$, $p$ and $\alpha$.
By simplifying the original system, we employ the duality approach used in [TZ09, Section 7], but following the spirit of [LIU14], where we avoid the need to prove stochastic Carleman estimates and use instead known results in the deterministic setting. Our result extends those in [BRT03] and [TZ09] to a wider range of values of $p$, circumventing technical challenges typically associated with stochastic control systems.
In this spirit, our final result brings together both applications and aims to establish a controllability result for the following controlled backward stochastic semilinear heat equation
$\begin{cases}\textnormal{d}{y}=\left(-\Delta y+f(y)+\chi_{\mathcal{O}_{0}}h\right)\textnormal{d}{t}+Y\textnormal{d}{W}(t)&\text{in }(0,T)\times\mathcal{O},\\ y=0&\text{on }(0,T)\times\partial\mathcal{O},\\ y(T)=y_{T}&\text{in }\mathcal{O}.\end{cases}$ (11)
In the $L^{2}$-setting, the (global) controllability of (11) in the sense of ˜1.8 has been studied in [HLP23] (see also [ZXL24]), under the assumption that $f$ is globally Lipschitz. However, as in the well-posedness results discussed earlier, this imposes a restrictive condition on the growth of $f$, which limits the applicability of the result.
A natural way to go beyond this limitation is to weaken the notion of controllability. In this direction, and in order to accommodate nonlinearities with weaker growth assumptions, we adopt the following notion of local controllability, which will serve as the framework for our result.
Definition 1.10 (Local controllability).
We say that equation (11) is locally controllable in $L^{\infty}$ if there exists $\delta>0$ such that for every terminal condition $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))$ with $\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\leq\delta$, there exists a control $h\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}_{0}))$ such that the weak solution $(y,Y)$ of (11) satisfies $y(0)=0$ a.s.
Using the control result in the linear setting stated in Proposition˜1.9 for $p=+\infty$ and in view of Theorem˜1.7, we can state the following result.
Theorem 1.11.
There is $\delta>0$ depending only on $T$, $\mathcal{O}$, and $\mathcal{O}_{0}$ such that for every $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))$ satisfying $\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\leq\delta$, there exists a control $h\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}_{0}))$ such that the weak solution $(y,Y)$ of (11) satisfies $y(0)=0$ a.s.
The proof of this result combines several techniques. First, we establish a null-controllability result for the linear backward equation with an $L^{\infty}$-weighted source term, crucially relying on the cost estimate (10). To achieve this, we adapt ideas from [LTT13] in the deterministic framework and from [HLP22] in the stochastic setting. For completeness, we present this intermediate result in our context here. The final step of the proof employs a Banach fixed-point argument, where arguments used in Theorem˜1.7 play a key role in controlling the nonlinear terms.
Remark 1.12.
Theorem 1.11 provides, to the best of our knowledge, the first true local controllability result for backward semilinear SPDEs. Previous works have not directly addressed this question. The closest related result appears in Section 4 of [HLP22], where the authors establish a weaker notion of controllability, named statisical local controllability: instead of reaching $y(0)=0$ a.s., they prove that $\mathbb{P}(y(0)=0)\geq 1-\varepsilon$ for some $\varepsilon>0$ sufficiently small starting from a small initial data. This distinction is significant, as our result guarantees local controllability in the almost sure sense.
It is also worth noting that [HLP22] considers the full backward equation, including the $\beta Y$ term (compare (9) and (11)). In contrast, our approach does not treat this term because of the limitations in the linear controllability analysis (see Proposition˜1.9). Nevertheless, it does establish a controllability result in a setting where previously only high-probability statements were known.
Remark 1.13.
Theorem˜1.11 does not cover the case in which the source term $f$ depends on $Y$. Indeed, in the key a priori estimate of the source-term method in Proposition˜5.1, the source term is required to lie in $L^{\infty}$, while our analysis only yields $L^{p}$-regularity for the process $Y$ (with finite $p$), which is generally insufficient to close the estimates. This obstruction already appears at the level of $L^{2}$, as noted in our previous work (see [HLP22, Remark 4.4]), where the source-term method does not provide suitable bounds when $f$ depends on $Y$. Addressing such nonlinear dependencies would likely require methods of a different nature, and we view this as an interesting direction for future research.
Remark 1.14.
For forward semilinear SPDEs of the form
$\begin{cases}\textnormal{d}y=(\Delta y+\chi_{\mathcal{O}_{0}}h+f(y))\,\textnormal{d}t+(ay+g(y))\,\textnormal{d}W(t)&\text{in }(0,T)\times\mathcal{O},\\ y=0&\text{on }(0,T)\times\partial\mathcal{O},\\ y(0)=y_{0}&\text{in }\mathcal{O},\end{cases}$ (12)
the basic controllability question is whether one can choose a suitable control $h$ (and possibly an additional control acting in the diffusion) to ensure
$y(T)=0\quad\text{a.s.}$
Let us recall that the linear case (i.e. $f\equiv 0$, $g\equiv 0$) has been studied extensively. Several controllability results are available under different assumptions, relying on Carleman estimates (see [FI96]) and on the Lebeau–Robbiano strategy [LR95], with both approaches requiring suitable adaptations to the stochastic setting (see, e.g., [BRT03, TZ09, LÜ11, LIU14]). We also point out that, in some of these works, an additional control acting in the diffusion term is required; namely, the term $ay\,\textnormal{d}W(t)$ in (12) is replaced by $(ay+H)\,\textnormal{d}W(t)$ for a suitably chosen control $H$.
When nonlinearities are present, particularly in the diffusion term, the analysis becomes substantially more delicate. In the globally Lipschitz setting, our previous work [HLP23] (see also [ZXL24]) established global controllability but—as in the linear case—only by introducing an additional control in the diffusion term. Subsequently, in [HLP22], we proved statistical local controllability with a single control (that is, $\mathbb{P}(y(T)=0)\geq 1-\varepsilon$ for $\varepsilon>0$ small and small initial data), under suitable structural assumptions on the nonlinear diffusion term $g(y)$, essentially in one spatial dimension. However, the general problem of proving true local controllability for semilinear equations with nonlinear noise remains open.
The present paper suggests a possible direction for improving forward results: the high-regularity estimates we obtain here for the backward equation appear to be precisely the type of regularity input that is currently missing in the forward analysis. Developing analogous high-regularity estimates for forward equations could plausibly lead to sharper controllability results, but this remains unexplored.
Remark 1.15.
The $L^{p}$- and $L^{\infty}$-estimates obtained in this work might also be relevant to stochastic optimal control for parabolic SPDEs. Backward equations naturally arise in Pontryagin-type maximum principles, and high-integrability or boundedness of the adjoint variables might play an important role in problems with bounded controls or state constraints. Although we do not pursue this direction here, the regularity tools developed in the present paper could extend to the optimal control setting. We refer the interested reader to [FHT12, DM13, LZ14, FL20] and the references therein for further developments in this direction. We hope that the present results may provide a useful starting point for future work, and we leave this as an interesting avenue for further investigation.
1.3 Outline of the paper
The rest of the paper is organized as follows. In Section˜2, we present and prove an Itô formula for the $L^{p}$-norm of the solution of the BSPDE (1). As a consequence, we will deduce the proof of Theorem˜1.3. Then, in Section˜3 we use the Itô formula to show uniform $L^{p}$-estimates then Theorem˜1.4 concerning $L^{\infty}$-estimates for the solution of the BSPDE (1). Section˜4 focuses on the study of the semilinear equation (7) and its well-posedness, while Section˜5 is devoted to prove the control results in Proposition˜1.9 and Theorem˜1.11. Lastly, in Section˜6 we give further remarks on the regularity obtained in Theorems 1.3 and 1.4.
2 Itô’s formula for the $L^{p}$-norm
The goal of this section is to state Itô’s formula for the $L^{p}$-norm and deduce $L^{p}$-estimates for the weak solution $(y,Y)$ of (1). Note that throughout this section, the $L^{p}$-estimates are not uniform with respect to $p$. The final conclusion will be the proof of Theorem˜1.3.
We begin by proving the following result, which is essential for the proofs of Theorem˜1.3 and Theorem˜1.4. It provides an initial a priori estimate together with an Itô’s formula for the $L^{p}$-norm of the solution to the backward SPDE (1). More precisely, we have the following.
Proposition 2.1 (Ito’s formula for $L^{p}$-norm).
Let $(y,Y)$ be a weak solution to (1) and assume that $y_{T}\in L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))$ and $F\in L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))$ for some $p\in[2,+\infty)$. Then, there is a constant $C>0$ only depending on $\mathcal{O}$, $\alpha$, $\beta$, and $p$ such that
${\mathbb{E}}\left(\sup_{t\in[0,T]}\|y(t)\|^{p}_{L^{p}(\mathcal{O})}\right)+{\mathbb{E}}\left(\int_{0}^{T}\int_{\mathcal{O}}|y|^{p-2}|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}t\right)$
$+{\mathbb{E}}\left(\int_{0}^{T}\int_{\mathcal{O}}|y|^{p-2}Y^{2}\,\textnormal{d}x\textnormal{d}t\right)\leq Ce^{CT}{\mathbb{E}}\left(\int_{\mathcal{O}}|y_{T}|^{p}\,\textnormal{d}x+\int_{0}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}t\right).$ (13)
Moreover, for any $t\in[0,T]$, we have
$\int_{\mathcal{O}}|y(t)|^{p}\,\textnormal{d}x+p(p-1)\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}s+\frac{p(p-1)}{2}\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}|Y|^{2}\,\textnormal{d}x\textnormal{d}s$
$=\int_{\mathcal{O}}|y_{T}|^{p}\,\textnormal{d}x+p\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}y\left(\alpha y+\beta Y+F\right)\,\textnormal{d}x\textnormal{d}{s}-p\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}yY\textnormal{d}{x}dW(s),\quad\textnormal{a.s.}$ (14)
For the proof, we follow the methodology in [DG15] adapted to the backward case and for the reader’s convenience we present it in several steps.
2.1 Technical lemmas
We begin by presenting some auxiliary functions. For each $n\in\mathbb{N}$ and some $p\geq 2$, consider the twice continuously differentiable function $\phi_{n}:\mathbb{R}\to\mathbb{R}$ defined by
$\phi_{n}(r):=\begin{cases}|r|^{p}&\text{ if }|r|<n,\\ n^{p-2}\frac{p(p-1)}{2}\left(|r|-n\right)^{2}+pn^{p-1}\left(|r|-n\right)+n^{p}&\text{ if }|r|\geq n.\end{cases}$ (15)
It is not difficult to check that $\phi_{n}$ satisfy
$|\phi_{n}(r)|\leq M|r|^{2},\qquad|\phi_{n}^{\prime}(r)|\leq M|r|,\qquad|\phi^{\prime\prime}(r)|\leq M,\qquad r\in{\mathbb{R}},$ (16)
where $M>0$ depends only on $p$ and $n$. Furthermore, we have
$|\phi_{n}(r)|\leq N|r|^{p},\qquad|\phi_{n}^{\prime}(r)|\leq N|r|^{p-1},\qquad|\phi^{\prime\prime}(r)|\leq N|r|^{p-2},\qquad r\in{\mathbb{R}},$ (17)
for a positive constant $N$ only depending on $p$, and if $n\to\infty$, we have
$\phi_{n}(r)\to|r|^{p},\qquad\phi_{n}^{\prime}(r)\to p|r|^{p-2}r,\qquad\phi_{n}^{\prime\prime}(r)\to p(p-1)|r|^{p-2},\qquad r\in{\mathbb{R}}.$ (18)
Additionally, we can check that the following inequalities hold
$|r\phi_{n}^{\prime}(r)|$ $\leq p\phi_{n}(r),$ (19)
$|\phi_{n}^{\prime}(r)|^{2}$ $\leq 4p\phi_{n}^{\prime\prime}(r)\phi_{n}(r),$ (20)
$[\phi^{\prime\prime}(r)]^{p/(p-2)}$ $\leq\left[p(p-1)\right]^{p/(p-2)}\phi_{n}(r)\qquad\textnormal{for all }r\in{\mathbb{R}},$ (21)
and, by the definition of (15), we can see that
$\phi_{n}^{\prime\prime}(r)\geq 0\quad\textnormal{for all $r\in\mathbb{R}$}.$ (22)
We start with the following lemma.
Lemma 2.2.
Let $n\in\mathbb{N}$, $p\in[2+\infty)$, $t\in[0,T]$, and $(y,Y)$ be a weak solution to (1). Then
${\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\,\textnormal{d}W(s)\right)=0.$
Proof.
By Cauchy-Schwarz and Hölder inequalities together with (16) we see that
$\int_{t}^{T}\left(\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\right)^{2}\textnormal{d}s$ $\leq\int_{t}^{T}\|\phi_{n}^{\prime}(y)\|^{2}_{L^{2}(\mathcal{O})}\|Y\|^{2}_{L^{2}(\mathcal{O})}\textnormal{d}s$
$\leq M\int_{t}^{T}\|y\|^{2}_{L^{2}(\mathcal{O})}\|Y\|^{2}_{L^{2}(\mathcal{O})}\textnormal{d}s$
$\leq M\sup_{s\in(t,T]}\|y(s)\|^{2}_{L^{2}(\mathcal{O})}\int_{t}^{T}\|Y\|^{2}_{L^{2}(\mathcal{O})}\textnormal{d}s,\quad\text{a.s.}$
for any $t\in[0,T]$. Taking expectation and using again Cauchy-Schwarz’s inequality we get
${\mathbb{E}}\left(\left[\int_{t}^{T}\left(\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\right)^{2}\textnormal{d}s\right]^{1/2}\right)\leq M{\mathbb{E}}\left(\sup_{s\in(t,T]}\|y(s)\|^{2}_{L^{2}(\mathcal{O})}\right)^{1/2}{\mathbb{E}}\left(\int_{t}^{T}\|Y\|^{2}_{L^{2}(\mathcal{O})}\textnormal{d}s\right)^{1/2}$
whence
${\mathbb{E}}\left(\left[\int_{t}^{T}\left(\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\right)^{2}\textnormal{d}s\right]^{1/2}\right)$
$\leq\frac{M}{2}\left[{\mathbb{E}}\left(\sup_{s\in(t,T]}\|y(s)\|^{2}_{L^{2}(\mathcal{O})}\right)+{\mathbb{E}}\left(\int_{t}^{T}\|Y\|^{2}_{L^{2}(\mathcal{O})}\textnormal{d}s\right)\right]<+\infty.$
by the regularity estimate (4). Defining $H(s):=\int_{\mathcal{O}}\phi_{n}^{\prime}(y(s))Y(s)\,\textnormal{d}x$, we have proved that $H(\cdot)\in L^{2}(\Omega\times(0,T))$. Therefore, ${\mathbb{E}}\left(\int_{t}^{T}H(s)\textnormal{d}{W}(s)\right)={\mathbb{E}}\left(\int_{0}^{T}H(s)\textnormal{d}{W}(s)-\int_{0}^{t}H(s)\textnormal{d}{W}(s)\right)=0$. This ends the proof. ∎
We continue with the following identity.
Lemma 2.3.
Let $n\in\mathbb{N}$ and $(y,Y)$ be a weak solution to (1). For any $t\in[0,T]$, the following identity holds
$\int_{\mathcal{O}}$ $\phi_{n}(y(t))\,\textnormal{d}x+\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime\prime}_{n}(y)|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}s+\frac{1}{2}\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime}_{n}(y)Y^{2}\,\textnormal{d}x\textnormal{d}s$
$=\int_{\mathcal{O}}\phi_{n}(y(T))\,\textnormal{d}x+\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime}_{n}(y)(\alpha y+\beta Y+F)\,\textnormal{d}x\textnormal{d}s-\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\textnormal{d}W(s),\quad\text{ a.s.}$ (23)
Proof.
Since $\phi_{n}$ is a twice continuously differentiable function, we can apply Itô’s formula and use equation (1) to get
$\int_{\mathcal{O}}\phi_{n}(y(T))\,\textnormal{d}x$ $=\int_{\mathcal{O}}\phi_{n}(y(t))\,\textnormal{d}x-\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime}_{n}(y)\Delta y\,\textnormal{d}x\textnormal{d}{s}-\int_{t}^{T}\int_{\mathcal{O}}\left(\phi^{\prime}_{n}(y)F-\frac{1}{2}\phi^{{}^{\prime\prime}}_{n}(y)Y^{2}\right)\,\textnormal{d}x\textnormal{d}s$
$-\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)(\alpha y+\beta Y)\,\textnormal{d}x\textnormal{d}{s}+\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{{}^{\prime}}(y)Y\,\textnormal{d}x\textnormal{d}W(s),\quad\text{ a.s.}$
for any $t\in[0,T]$. Note that for $i=1,\ldots,d$, $\partial_{x_{i}}(\phi_{n}^{\prime}(y))=\phi_{n}^{\prime\prime}(y)\partial_{x_{i}}y$, whence integrating by parts in the space variable and using the homogeneous Dirichlet boundary condition of (1) yields
$\int_{\mathcal{O}}\phi_{n}(y(T))\,\textnormal{d}x$ $=\int_{\mathcal{O}}\phi_{n}(y(t))\,\textnormal{d}x+\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime\prime}_{n}(y)|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}s-\int_{t}^{T}\int_{\mathcal{O}}\left(\phi^{\prime}_{n}(y)F-\frac{1}{2}\phi^{\prime\prime}_{n}(y)Y^{2}\right)\,\textnormal{d}x\textnormal{d}s$
$-\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)(\alpha y+\beta Y)\,\textnormal{d}x\textnormal{d}{s}+\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\textnormal{d}W(s),\quad\text{ a.s.}$
After rearranging some terms, we obtain the desired equality. ∎
In the remainder of this section, $C$ denotes a generic positive constant uniform with respect to $n$ and that may change from line to line. We have the following estimate.
Lemma 2.4.
Let $n\in\mathbb{N}$, $p\in[2,+\infty)$, $(y,Y)$ be a weak solution to (1) and assume that $y_{T}\in L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))$ and $F\in L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))$. Then, the following estimate holds
${\mathbb{E}}$ $\left(\int_{0}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime\prime}(y)|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}s\right)+{\mathbb{E}}\left(\int_{0}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime\prime}(y)Y^{2}\,\textnormal{d}x\textnormal{d}s\right)$ (24)
$\leq Ce^{CT}{\mathbb{E}}\left(\int_{\mathcal{O}}|y_{T}|^{p}\,\textnormal{d}x+\int_{0}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}s\right),$
for some constant $C>0$ that only depends on $p$, $\|\alpha\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$ and $\|\beta\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$.
Proof.
Let $t>0$ and $0<r\leq t$. From identity (23) of Lemma˜2.3, we take conditional expectation to obtain
${\mathbb{E}}\left(\int_{\mathcal{O}}\right$. $\left.\phi_{n}(y(t))\,\textnormal{d}x+\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime\prime}_{n}(y)|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}s+\frac{1}{2}\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime\prime}_{n}(y)Y^{2}\,\textnormal{d}x\textnormal{d}s\;{\Big|}\;\mathcal{F}_{r}\right)$
$={\mathbb{E}}\left(\int_{\mathcal{O}}\phi_{n}(y(T))\,\textnormal{d}x+\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime}_{n}(y)(\alpha y+\beta Y+F)\,\textnormal{d}x\textnormal{d}s\;{\Big|}\;\mathcal{F}_{r}\right)$
where we have used that ${\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\textnormal{d}W(s)\mid\mathcal{F}_{r}\right)=0$ due to Lemma˜2.2. From (19)-(21), the positivity property (22) and Young’s inequality, we can deduce
$|\phi_{n}^{\prime}(y)F|$ $\leq C|F|\phi_{n}(y)^{(p-1)/p}\leq C\big(|F|^{p}+\phi_{n}(y)\big),$ (25)
$|\phi_{n}^{\prime}(y)\alpha y|$ $\leq C|\alpha|\phi_{n}(y),$ (26)
$|\phi_{n}^{\prime}(y)\beta Y|$ $\leq\delta\phi_{n}^{\prime\prime}(y)|Y|^{2}+\tfrac{C|\beta|^{2}}{\delta}\phi_{n}(y),$ (27)
almost surely, for a constant $C>0$ only depending on $p$ and any $\delta\in(0,1)$. Therefore, taking $\delta$ small enough
${\mathbb{E}}\left(\int_{\mathcal{O}}\right$. $\left.\phi_{n}(y(t))\,\textnormal{d}x+\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime\prime}_{n}(y)|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}s+\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime\prime}_{n}(y)Y^{2}\,\textnormal{d}x\textnormal{d}s\;{\Big|}\;\mathcal{F}_{r}\right)$
$\leq C{\mathbb{E}}\left(\int_{\mathcal{O}}\phi_{n}(y(T))\,\textnormal{d}x+\int_{t}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}s+\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}(y)\,\textnormal{d}x\textnormal{d}s\;{\Big|}\;\mathcal{F}_{r}\right),$
where $C>0$ depends on $p$ and the norms $\|\alpha\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$ and $\|\beta\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$.
Taking expectation in the above inequality and by Fubini’s theorem and Lemma˜2.2, we get
${\mathbb{E}}\left(\int_{\mathcal{O}}\right$. $\left.\phi_{n}(y(t))\,\textnormal{d}x+\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime\prime}_{n}(y)|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}s+\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime\prime}_{n}(y)Y^{2}\,\textnormal{d}x\textnormal{d}s\right)$
$\leq C{\mathbb{E}}\left(\int_{\mathcal{O}}\phi_{n}(y(T))\,\textnormal{d}x+\int_{0}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}s\right)+C\int_{t}^{T}{\mathbb{E}}\left(\int_{\mathcal{O}}\phi_{n}(y)\,\textnormal{d}x\textnormal{d}s\right).$ (28)
By (22), we can drop the second and third terms in the above estimate and, by the backward Grönwall inequality (see Lemma˜A.1), we obtain
${\mathbb{E}}\left(\int_{\mathcal{O}}\phi_{n}(y(t))\,\textnormal{d}x\right)\leq Ce^{CT}{\mathbb{E}}\left(\int_{\mathcal{O}}\phi_{n}(y(T))\,\textnormal{d}x+\int_{0}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}s\right).$ (29)
Integrating in time (29) and combining with (28) entails
${\mathbb{E}}$ $\left(\int_{\mathcal{O}}\phi_{n}(y(t))\,\textnormal{d}x\right)+{\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime\prime}(y)|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}s\right)+{\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime\prime}(y)Y^{2}\,\textnormal{d}x\textnormal{d}s\right)$
$\leq Ce^{CT}\left[{\mathbb{E}}\left(\int_{\mathcal{O}}\phi_{n}(y(T))\,\textnormal{d}x\right)+{\mathbb{E}}\left(\int_{0}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}s\right)\right],$ (30)
where we have replaced the terminal data of (1). Dropping the first term in (30) and applying the monotone convergence theorem yield the desired result. ∎
The following result tells us that we can improve Lemma˜2.4 by adding a supremum (in time) estimate on the left-hand side of (24).
Lemma 2.5.
Under the assumptions of Lemma˜2.4, there is $C>0$ only depending on $p$ such that
${\mathbb{E}}\left(\sup_{t\in[0,T]}\int_{\mathcal{O}}\phi_{n}(y(t))\right)+{\mathbb{E}}\left(\int_{0}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime\prime}(y)|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}s\right)+{\mathbb{E}}\left(\int_{0}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime\prime}(y)Y^{2}\,\textnormal{d}x\textnormal{d}s\right)$ (31)
$\quad\leq Ce^{CT}{\mathbb{E}}\left(\int_{\mathcal{O}}|y_{T}|^{p}\,\textnormal{d}x+\int_{0}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}s\right).$
Proof.
Using as a starting point identity (23) of Lemma˜2.3, we use (25)–(27) to deduce that almost surely
$\int_{\mathcal{O}}$ $\phi_{n}(y(t))\,\textnormal{d}x+\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime\prime}_{n}(y)|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}s+\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime\prime}_{n}(y)Y^{2}\,\textnormal{d}x\textnormal{d}{s}$
$\leq C\left(\int_{\mathcal{O}}|y_{T}|^{p}\,\textnormal{d}x+\int_{t}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}s+\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}(y)\,\textnormal{d}x\textnormal{d}s\right)-\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\textnormal{d}W(s),$
for any $t\in[0,T]$ and some $C>0$ only depending on $p$, $\|\alpha\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$ and $\|\beta\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$. Hence
$\int_{\mathcal{O}}\phi_{n}(y(t))\,\textnormal{d}x\leq C\left(\int_{\mathcal{O}}|y_{T}|^{p}\,\textnormal{d}x+\int_{0}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}s\right)+\sup_{t\in[0,T]}\left|\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\textnormal{d}W(s)\right|$
$+C\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}(y)\,\textnormal{d}x\textnormal{d}{s},\quad\text{ a.s.}$ (32)
From the backward Grönwall inequality (see Lemma˜A.1) and taking supremum in $[0,T]$, we have that a.s.
$\sup_{t\in[0,T]}$ $\int_{\mathcal{O}}\phi_{n}(y(t))\,\textnormal{d}x$
$\leq Ce^{CT}\left(\int_{\mathcal{O}}|y_{T}|^{p}\,\textnormal{d}x+\int_{0}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}s+2\sup_{t\in[0,T]}\left|\int_{0}^{t}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}xdW(s)\right|\right),$ (33)
where we have used that
$\sup_{t\in[0,T]}$ $\left|\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}xdW(s)\right|$
$\leq\left|\int_{0}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\textnormal{d}{W}(s)\right|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\textnormal{d}{W}(s)\right|$
$\leq 2\sup_{t\in[0,T]}\left|\int_{0}^{t}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\textnormal{d}{W}(s)\right|.$ (34)
We take expectation in (33) and apply Burkholder-Davis-Gundy inequality to get
${\mathbb{E}}\left(\sup_{t\in[0,T]}\int_{\mathcal{O}}\phi_{n}(y(t))\,\textnormal{d}x\right)$
$\leq Ce^{CT}\left[{\mathbb{E}}\left(\int_{\mathcal{O}}|y_{T}|^{p}\,\textnormal{d}x+\int_{0}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}s\right)+{\mathbb{E}}\left(\left|\int_{0}^{T}\left(\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\right)^{2}\textnormal{d}s\right|^{1/2}\right)\right].$ (35)
To estimate the last term in the above inequality we note that from (20)
$|\phi_{n}^{\prime}(y)Y|\leq 2|Y|\phi_{n}^{\prime\prime}(y)^{1/2}\phi_{n}(y)^{1/2}\quad\text{a.s.}$
and applying Hölder and Young inequalities we have that there exists a numerical constant $C>0$ such that
${\mathbb{E}}\left(\left|\int_{0}^{T}\left(\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\right)^{2}\textnormal{d}s\right|^{1/2}\right)$
$\leq\delta{\mathbb{E}}\left(\sup_{s\in[0,T]}\int_{\mathcal{O}}\phi_{n}(y)\,\textnormal{d}x\right)+\frac{C}{\delta}{\mathbb{E}}\left(\int_{0}^{T}\int_{\mathcal{O}}|Y|^{2}|\phi_{n}^{\prime\prime}(y)|\,\textnormal{d}x\textnormal{d}s\right),$ (36)
for any $\delta\in(0,1)$. Replacing (36) in (35) and taking $\delta$ small enough we get
${\mathbb{E}}$ $\left(\sup_{t\in[0,T]}\int_{\mathcal{O}}\phi_{n}(y(t))\,\textnormal{d}x\right)$
$\leq Ce^{CT}{\mathbb{E}}\left(\int_{\mathcal{O}}|y_{T}|^{p}\,\textnormal{d}x+\int_{0}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}s+\int_{0}^{T}\int_{\mathcal{O}}|Y|^{2}|\phi_{n}^{\prime\prime}(y)|\,\textnormal{d}x\textnormal{d}s\right).$ (37)
Finally, putting together estimate (24) of Lemma 2.4 and inequality (37) we obtain the desired result. This ends the proof. ∎
2.2 Proof of the Ito’s formula for the $L^{p}$-norm
Now we are in position to prove the Ito’s formula, that is, Proposition˜2.1.
Proof of Proposition˜2.1.
Estimate (13), follows from (18), estimate (31) in Lemma˜2.5 and a straightfoward application of Fatou’s Lemma. This proves the first assertion of Theorem˜1.3.
To prove the Itô’s formula (14), we proceed as follows. From Lemma˜2.3, we have that the following identity holds
$\int_{\mathcal{O}}$ $\phi_{n}(y(t))\,\textnormal{d}x+\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime\prime}_{n}(y)|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}s+\frac{1}{2}\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime\prime}_{n}(y)Y^{2}\,\textnormal{d}x\textnormal{d}s$
$=\int_{\mathcal{O}}\phi_{n}(y(T))\,\textnormal{d}x+\int_{t}^{T}\int_{\mathcal{O}}\phi^{\prime}_{n}(y)(\alpha y+\beta Y+F)\,\textnormal{d}x\textnormal{d}s-\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\textnormal{d}W(s)\quad\text{ a.s.,}$ (38)
for all $t\in[0,T]$. We will see that each term in (38) converges to the corresponding one in (14) as $n\to\infty$. At this point, we will make use of estimate (13).
From estimate (13), we readily deduce that
$\sup_{t\in[0,T]}\|y(t)\|^{p}_{L^{p}(\mathcal{O})}<+\infty,\quad\textnormal{a.s.}$ (39)
Then, by (17), (18), (39), and dominated convergence theorem we obtain that
$\int_{\mathcal{O}}\phi_{n}(y(t))\,\textnormal{d}x$ $\to\int_{\mathcal{O}}|y(t)|^{p}\,\textnormal{d}x\quad\text{ a.s. for all $t\in[0,T]$}.$
This proves the convergence of the first terms in both sides of identity (38).
Now we note from (17) that a.s.
$|\phi_{n}^{\prime}(y)F|\leq N|y|^{p-1}|F|$
and, moreover, by succesive applications of Hölder inequality we have for all $t\in[0,T]$
$\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-1}|F|\,\textnormal{d}x\textnormal{d}s$ $\leq\|y\|_{L^{p}((t,T];L^{p}(\mathcal{O}))}^{p-1}\|F\|_{L^{p}((t,T];L^{p}(\mathcal{O}))}<+\infty\quad\text{ a.s.},$
where the boundedness follows from (39) and the regularity of $F$. Therefore, dominated convergence theorem together with (18) imply that
$\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)F\,\textnormal{d}x\textnormal{d}s\to p\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}yF\,\textnormal{d}x\textnormal{d}s\quad\text{ a.s. for all $t\in[0,T]$.}$
Similar arguments using the energy estimate (13) yield that
$\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime\prime}(y)|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}s$ $\to p(p-1)\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}{s}\quad\text{ a.s.,}$
$\frac{1}{2}\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime\prime}(y)Y^{2}\,\textnormal{d}x\textnormal{d}s$ $\to\frac{p(p-1)}{2}\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}|Y|^{2}\,\textnormal{d}x\textnormal{d}{s}\quad\text{ a.s.,}$
$\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)\alpha y\,\textnormal{d}x\textnormal{d}s$ $\to p\int_{t}^{T}\int_{\mathcal{O}}\alpha|y|^{p}\,\textnormal{d}x\textnormal{d}{s}\quad\text{ a.s.,}$
$\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)\beta Y\,\textnormal{d}x\textnormal{d}s$ $\to p\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}\beta yY\,\textnormal{d}x\textnormal{d}{s}\quad\text{ a.s.,}$
for all $t\in[0,T]$.
Finally, for the term $\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\,\textnormal{d}x\textnormal{d}W(s)$, first we will prove that
$\int_{t}^{T}\left|\langle\phi_{n}^{\prime}(y),Y\rangle_{L^{2}(\mathcal{O})}-\left\langle p|y|^{p-2}y,Y\right\rangle_{L^{2}(\mathcal{O})}\right|^{2}\textnormal{d}s\to 0\quad\text{ a.s. for all $t\in[0,T]$}.$ (40)
For this, we recall from (17) that $|\phi^{\prime}_{n}(r)|\leq N|r|^{p-1}$ and by Cauchy-Schwarz inequality
$\left|\langle\phi_{n}^{\prime}(y),Y\rangle_{L^{2}(\mathcal{O})}-\langle p|y|^{p-2}y,Y\rangle_{L^{2}(\mathcal{O})}\right|^{2}$ $\leq C\left(\int_{\mathcal{O}}|y|^{p-2}|Y|^{2}\,\textnormal{d}x\right)\left(\int_{\mathcal{O}}|y|^{p}\,\textnormal{d}x\right)\quad\textnormal{a.s.}$
On the other hand, by estimate (13), we have
$\int_{t}^{T}$ $\left(\int_{\mathcal{O}}|y|^{p-2}|Y|^{2}\,\textnormal{d}x\right)\left(\int_{\mathcal{O}}|y|^{p}\,\textnormal{d}x\right)\textnormal{d}{s}$
$\leq C\left(\sup_{s\in[t,T]}\int_{\mathcal{O}}|y(s)|^{p}\,\textnormal{d}x\right)\left(\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}|Y|^{2}\,\textnormal{d}x\textnormal{d}{s}\right),\quad$
<+, $<+\infty,\quad\text{ a.s. for all $t\in[0,T]$.}$
From (18), the continuity of the $L^{2}$-inner product, and dominated convergence theorem allow us to conclude (40) and therefore
$\int_{t}^{T}\int_{\mathcal{O}}\phi_{n}^{\prime}(y)Y\textnormal{d}W(s)\to p\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}yY\textnormal{d}W(s)\;\;\text{ a.s. for all $t\in[0,T]$.}$
The proof is complete. ∎
2.3 Extra regularity for the process $Y$ and proof of Theorem˜1.3
The goal of this part is to establish extra regularity for the process $Y$.
Proposition 2.6.
Let $(y,Y)$ be a weak solution to (1) and assume that $y_{T}\in L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))$ and $F\in L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))$ for some $p\in(2,+\infty)$. Then $Y\in L^{p}_{\mathbb{F}}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O})))$ and there is a constant $C>0$ only depending on $\mathcal{O}$, $\alpha$, $\beta$, and $p$ such that
${\mathbb{E}}\left[\left(\int_{0}^{T}\|Y(t)\|_{L^{2}(\mathcal{O})}^{2}\textnormal{d}t\right)^{p/2}\right]\leq Ce^{CT}{\mathbb{E}}\left(\|y_{T}\|^{p}_{L^{p}(\mathcal{O})}+\int_{0}^{T}\|F(t)\|_{L^{p}(\mathcal{O})}^{p}\textnormal{d}t\right).$ (41)
Proof.
Let $(y,Y)$ be a weak solution to (1). Along the proof, $C$ stands for a generic positive constant depending at most on $\mathcal{O}$, $\|\alpha\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$, $\|\beta\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$, and $p$ but which is uniform with respect to $T$. This constant may change from line to line.
Using formula (14) in Proposition˜2.1 with $p=2$ (which amounts to the usual Ito’s formula) and dropping some of the positive terms we have
$\int_{t}^{T}\int_{\mathcal{O}}|Y|^{2}\,\textnormal{d}x\textnormal{d}{s}\leq\int_{\mathcal{O}}|y_{T}|^{2}\,\textnormal{d}x+2\int_{t}^{T}\int_{\mathcal{O}}y(\alpha y+\beta Y+F)\,\textnormal{d}x\textnormal{d}{s}-2\int_{t}^{T}\int_{\mathcal{O}}yY\,\textnormal{d}x\textnormal{d}{W}(s),\quad\textnormal{a.s.}$
for any $t\in[0,T]$. Since $y\in L^{2}_{\mathbb{F}}(\Omega;C([0,T];L^{2}(\mathcal{O})))$, then $\int_{\mathcal{O}}|y_{T}|^{2}\leq\sup_{s\in[0,T]}\|y(s)\|^{2}_{L^{2}(\mathcal{O})}$ a.s. From this, and taking the power $p/2$ with $p>2$ in both sides of the above inequality, we deduce
$\left(\int_{t}^{T}\int_{\mathcal{O}}|Y|^{2}\,\textnormal{d}x\textnormal{d}{s}\right)^{p/2}$
$\;\leq C\left(\sup_{s\in[0,T]}\|y(s)\|_{L^{2}(\mathcal{O})}^{p}+\left(\int_{0}^{T}\int_{\mathcal{O}}|\alpha||y|^{2}\,\textnormal{d}x\textnormal{d}{s}\right)^{p/2}+\left(\int_{0}^{T}\int_{\mathcal{O}}|\beta yY|\,\textnormal{d}x\textnormal{d}{s}\right)^{p/2}\right)$
$\quad+C\left(\left(\int_{0}^{T}\int_{\mathcal{O}}|yF|\,\textnormal{d}x\textnormal{d}{s}\right)^{p/2}+\left|\int_{t}^{T}\int_{\mathcal{O}}yY\,\textnormal{d}x\textnormal{d}{W}(s)\right|^{p/2}\right),$ (42)
almost surely for any $t\in[0,T]$.
Let us estimate the expectation of the last four terms in the right-hand side of (42). For the first one, from Hölder inequality and since $\alpha$ verifies (H), we readily deduce
${\mathbb{E}}\left[\left(\int_{0}^{T}\int_{\mathcal{O}}|\alpha||y|^{2}\,\textnormal{d}x\textnormal{d}{s}\right)^{p/2}\right]\leq CT^{p/2}{\mathbb{E}}\left(\sup_{s\in[0,T]}\|y(s)\|_{L^{2}(\mathcal{O})}^{p}\right).$ (43)
For the second one, recalling (H) and using Cauchy-Schwarz inequality successively, we have
$\left(\int_{0}^{T}\int_{\mathcal{O}}|\beta yY|\,\textnormal{d}x\textnormal{d}{s}\right)^{p/2}$ $\leq C\left(\int_{0}^{T}\|y(s)\|^{2}_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right)^{p/4}\left(\int_{0}^{T}\|Y(s)\|^{2}_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right)^{p/4}\quad\textnormal{a.s.}$
Therefore, taking expectation and using Cauchy-Schwarz and Young inequalities we obtain
${\mathbb{E}}\left[\left(\int_{0}^{T}\int_{\mathcal{O}}|\beta yY|\,\textnormal{d}x\textnormal{d}{s}\right)^{p/2}\right]$
$\qquad\leq\delta{\mathbb{E}}\left[\left(\int_{0}^{T}\|Y(s)\|^{2}_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right)^{p/2}\right]+\frac{C}{\delta}{\mathbb{E}}\left[\left(\int_{0}^{T}\|y(s)\|^{2}_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right)^{p/2}\right]$
for any $\delta>0$. Using Hölder inequality in the last term, we can further estimate
${\mathbb{E}}\left[\left(\int_{0}^{T}\int_{\mathcal{O}}|\beta yY|\,\textnormal{d}x\textnormal{d}{s}\right)^{p/2}\right]$
[…]
To estimate the third term in the right-hand side of (42), we proceed as follows. Using Cauchy-Schwarz and Hölder inequalities, we get
$\left(\int_{0}^{T}\int_{\mathcal{O}}|yF|\,\textnormal{d}x\textnormal{d}{s}\right)^{p/2}\leq\sup_{s\in[0,T]}\|y(s)\|^{p/2}_{L^{2}(\mathcal{O})}\left(\int_{0}^{T}\|F(s)\|_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right)^{p/2},\quad\textnormal{a.s.}$
Taking expectation and using Cauchy-Schwarz and Young inequalities we obtain
${\mathbb{E}}\left[\left(\int_{0}^{T}\int_{\mathcal{O}}|yF|\,\textnormal{d}x\textnormal{d}{s}\right)^{p/2}\right]\leq\frac{1}{2}{\mathbb{E}}\left(\sup_{s\in[0,T]}\|y(s)\|^{p}_{L^{2}(\mathcal{O})}\right)+\frac{1}{2}{\mathbb{E}}\left[\left(\int_{0}^{T}\|F(s)\|_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right)^{p}\right].$ (45)
To estimate the last term in (42), arguing as we did in (34), we have that
$\sup_{t\in[0,T]}\left|\int_{t}^{T}\int_{\mathcal{O}}yY\,\textnormal{d}x\textnormal{d}{W}(s)\right|^{p/2}\leq 2\sup_{t\in[0,T]}\left|\int_{0}^{t}\int_{\mathcal{O}}yY\,\textnormal{d}x\textnormal{d}{W}(s)\right|^{p/2},$
whence, by Burkholder-Davis-Gundy inequality, we deduce that
${\mathbb{E}}\left[\left|\int_{t}^{T}\int_{\mathcal{O}}yY\,\textnormal{d}x\textnormal{d}{W}(s)\right|^{p/2}\right]\leq C{\mathbb{E}}\left[\left|\int_{0}^{T}\left(\int_{\mathcal{O}}yY\,\textnormal{d}x\right)^{2}\textnormal{d}{s}\right|^{p/4}\right].$ (46)
Using Cauchy-Schwarz and Hölder inequalities we can estimate the right-hand side of (46) as
${\mathbb{E}}\left[\left|\int_{t}^{T}\int_{\mathcal{O}}yY\,\textnormal{d}x\textnormal{d}{W}(s)\right|^{p/2}\right]$ $\leq C{\mathbb{E}}\left[\sup_{s\in[0,T]}\|y(s)\|_{L^{2}(\mathcal{O})}^{p/2}\left|\int_{0}^{T}\|Y(s)\|^{2}_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right|^{p/4}\right],$
and from Young inequality, we deduce that for any $\delta>0$
${\mathbb{E}}\left[\left|\int_{t}^{T}\int_{\mathcal{O}}yY\,\textnormal{d}x\textnormal{d}{W}(s)\right|^{p/2}\right]$ $\leq\delta{\mathbb{E}}\left[\left|\int_{0}^{T}\|Y(s)\|^{2}_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right|^{p/2}\right]+\frac{C}{\delta}{\mathbb{E}}\left(\sup_{s\in[0,T]}\|y(s)\|_{L^{2}(\mathcal{O})}^{p}\right).$ (47)
Taking expectation in (42) and putting together with (43), (44), (45), and (47) we have that for any $t\in[0,T]$ the following holds
${\mathbb{E}}\left[\left(\int_{t}^{T}\int_{\mathcal{O}}|Y|^{2}\,\textnormal{d}x\textnormal{d}{s}\right)^{p/2}\right]$
$\quad\leq C\left(1+\frac{2T^{p/2}}{\delta}\right){\mathbb{E}}\left(\sup_{s\in[0,T]}\|y(s)\|_{L^{2}(\mathcal{O})}^{p}\right)+2\delta{\mathbb{E}}\left[\left(\int_{0}^{T}\|Y(s)\|^{2}_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right)^{p/2}\right]$
$\qquad+\frac{1}{2}{\mathbb{E}}\left[\left(\int_{0}^{T}\|F(s)\|_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right)^{p}\right].$
In particular, taking $t=0$ in the above estimate and setting $\delta>0$ small enough yields
${\mathbb{E}}\left[\left(\int_{0}^{T}\|Y(s)\|^{2}_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right)^{p/2}\right]$
$\quad\leq C(1+T^{p/2}){\mathbb{E}}\left(\sup_{s\in[0,T]}\|y(s)\|_{L^{2}(\mathcal{O})}^{p}\right)+C{\mathbb{E}}\left[\left(\int_{0}^{T}\|F(s)\|_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right)^{p}\right].$ (48)
To conclude, using Hölder inequality and since $\|z\|^{p}_{L^{2}(\mathcal{O})}\leq|\mathcal{O}|^{\frac{p-2}{2}}\|z\|^{p}_{L^{p}(\mathcal{O})}$ for any $z\in L^{p}(\mathcal{O})$, we obtain
${\mathbb{E}}\left[\left(\int_{0}^{T}\|Y(s)\|_{L^{2}(\mathcal{O})}^{2}\textnormal{d}{s}\right)^{p/2}\right]$
$\quad\leq C(1+T^{p/2}){\mathbb{E}}\left(\sup_{s\in[0,T]}\|y(s)\|_{L^{2}(\mathcal{O})}^{p}\right)+CT^{p-1}{\mathbb{E}}\left(\int_{0}^{T}\|F(s)\|_{L^{2}(\mathcal{O})}^{p}\textnormal{d}{s}\right)$
$\quad\leq C^{\prime}e^{C^{\prime}T}{\mathbb{E}}\left(\sup_{s\in[0,T]}\|y(s)\|_{L^{p}(\mathcal{O})}^{p}+\int_{0}^{T}\|F(s)\|_{L^{p}(\mathcal{O})}^{p}\textnormal{d}{s}\right),$ (49)
for some constant $C^{\prime}>0$ only depending on $\mathcal{O}$, $\alpha$, $\beta$, and $p$. To conclude, we use estimate (13) in (49) to obtain the desired inequality (41). This ends the proof. ∎
Proof of Theorem˜1.3.
The result follows directly from Proposition˜2.1 and Proposition˜2.6. ∎
3 Uniform $L^{p}$- and $L^{\infty}$-estimates
The goal of this section is to show Theorem˜1.4. In the first step towards Theorem˜1.4, we will use Itô’s formula for the $L^{p}$-norm (see Proposition˜2.1) to obtain a precise uniform estimate for the solution to (1). The result reads as follows.
Proposition 3.1.
Let $(y,Y)$ be a weak solution to (1). Assume that $y_{T}\in L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))$ and $F\in L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))$. Then the process $y$ taken from the solution $(y,Y)$ to (1) satisfies
$\left\lVert y\right\rVert_{L^{\infty}_{\mathbb{F}}(0,T;L^{p}(\Omega;L^{p}(\mathcal{O})))}\leq\exp(CT)\left(\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))}+\left\lVert F\right\rVert_{L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))}\right),$
for a constant $C>0$ independent of $y_{T}$, $F$ and $p$.
Proof.
Let $(y,Y)$ be a weak solution to (1). We start from formula (14) in Proposition˜2.1, namely
$\int_{\mathcal{O}}|y(t)|^{p}\,\textnormal{d}x+p(p-1)\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}|\nabla y|^{2}\,\textnormal{d}x\textnormal{d}s+\frac{p(p-1)}{2}\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}|Y|^{2}\,\textnormal{d}x\textnormal{d}s$
$=\int_{\mathcal{O}}|y_{T}|^{p}\,\textnormal{d}x+p\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}y\left(\alpha y+\beta Y+F\right)\,\textnormal{d}x-p\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}yYdW(s),\quad\textnormal{a.s.,}$
that holds for any $t\in[0,T]$.
In particular, by taking into account that the third term in the left-hand side is nonnegative, we deduce that
$\int_{\mathcal{O}}|y(t)|^{p}\,\textnormal{d}x+\frac{p(p-1)}{2}\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}|Y|^{2}\,\textnormal{d}x\textnormal{d}s$
$\quad\leq\int_{\mathcal{O}}|y_{T}|^{p}\,\textnormal{d}x+p\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-1}\left|\alpha y+\beta Y+F\right|\,\textnormal{d}x-p\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}yY\textnormal{d}W(s),\quad\text{a.s.}$ (50)
Let $r\leq t$ and take conditional expectation with respect to $\mathcal{F}_{r}$ in (50) to get
$\mathbb{E}\left(\left\lVert y(t)\right\rVert_{L^{p}(\mathcal{O})}^{p}\;{\Big|}\;\mathcal{F}_{r}\right)+\frac{p(p-1)}{2}{\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}|Y|^{2}\,\textnormal{d}x\textnormal{d}s\;{\Big|}\;\mathcal{F}_{r}\right)$
$\quad\leq\mathbb{E}\left(\left\lVert y_{T}\right\rVert_{L^{p}(\mathcal{O})}^{p}\;{\Big|}\;\mathcal{F}_{r}\right)+p\,\mathbb{E}\left(\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-1}\big[|\alpha y|+|\beta Y|+|F|\big]\,\textnormal{d}x\textnormal{d}{s}\;{\Big|}\;\mathcal{F}_{r}\right),$ (51)
because
$\mathbb{E}\left(\int_{t}^{T}\int_{\mathcal{O}}p|y|^{p-2}yYdW(s)\;{\Big|}\;\mathcal{F}_{r}\right)=0.$
This can be shown by employing estimate (13) and following the steps in Lemma˜2.2. For brevity, we skip the details.
Let us estimate the second integral in (51). By a direct computation using the monotonicty of the conditional expectation, we have
$p\,\mathbb{E}\left(\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-1}|\alpha y|\,\textnormal{d}x\textnormal{d}{s}\;{\Big|}\;\mathcal{F}_{r}\right)\leq p\,\|\alpha\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\,\mathbb{E}\left(\int_{t}^{T}\int_{\mathcal{O}}|y|^{p}\,\textnormal{d}x\textnormal{d}{s}\;{\Big|}\;\mathcal{F}_{r}\right).$ (52)
For the second one, by Cauchy-Schwarz and Young inequality we have
$p\,\mathbb{E}\left(\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-1}|\beta Y|\,\textnormal{d}x\textnormal{d}{s}\;{\Big|}\;\mathcal{F}_{r}\right)$
$\quad\leq\frac{p}{2}\,{\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}|Y|^{2}\,\textnormal{d}x\textnormal{d}{s}\;{\Big|}\;\mathcal{F}_{r}\right)+\frac{p\|\beta\|^{2}_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}}{2}{\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}|y|^{p}\,\textnormal{d}x\textnormal{d}{s}\;{\Big|}\;\mathcal{F}_{r}\right).$ (53)
For the last one, using Young inequality $ab\leq(1/p)a^{p}+(1/q)b^{q}$ with q=p/(p1)q=p/(p-1), so $pab\leq a^{p}+(p-1)b^{p/(p-1)}$ with $a=|F|$ and $b=|y|^{p-1}$ we obtain that
$p\,\mathbb{E}\left(\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-1}|F|\,\textnormal{d}x\textnormal{d}{s}\;{\Big|}\;\mathcal{F}_{r}\right)$
$\leq{\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}{s}\;{\Big|}\;\mathcal{F}_{r}\right)+(p-1){\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}|y|^{p}\,\textnormal{d}x\textnormal{d}{s}|\mathcal{F}_{r}\right).$ (54)
Therefore, combining (51) with estimates (52)–(54) we get
$\mathbb{E}\left(\left\lVert y(t)\right\rVert_{L^{p}(\mathcal{O})}^{p}\;{\Big|}\;\mathcal{F}_{r}\right)+\frac{p(p-2)}{2}{\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}|Y|^{2}\,\textnormal{d}x\textnormal{d}s\;{\Big|}\;\mathcal{F}_{r}\right)$
$\quad\leq\mathbb{E}\left(\left\lVert y_{T}\right\rVert_{L^{p}(\mathcal{O})}^{p}\;{\Big|}\;\mathcal{F}_{r}\right)+{\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}{s}\;{\Big|}\;\mathcal{F}_{r}\right)$
$\qquad+[pK+(p-1)]\,\mathbb{E}\left(\int_{t}^{T}\int_{\mathcal{O}}|y|^{p}\,\textnormal{d}x\textnormal{d}{s}\;{\Big|}\;\mathcal{F}_{r}\right),$
where $K:=\|\alpha\|_{L^{\infty}_{\mathbb{F}}(0,T,L^{\infty}(\mathcal{O}))}+\|\beta\|^{2}_{L^{\infty}_{\mathbb{F}}(0,T,L^{\infty}(\mathcal{O}))}$. Note that since $p\geq 2$, the second term in the left-hand side of the above inequality is always nonnegative.
Now, we take $r=t$ in the previous estimate and use the fact that $\mathbb{E}(\left\lVert y(t)\right\rVert_{L^{p}(\mathcal{O})}^{p}\mid\mathcal{F}_{t})=\left\lVert y(t)\right\rVert_{L^{p}(\mathcal{O})}^{p}$ since $y$ is $\mathcal{F}_{t}$-adapted to obtain
$\left\lVert y(t)\right\rVert_{L^{p}(\mathcal{O})}^{p}$
$\leq{\mathbb{E}}\left(\left\lVert y_{T}\right\rVert_{L^{p}(\mathcal{O})}^{p}\;{\Big|}\;\mathcal{F}_{t}\right)+{\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}|F|^{p}\,\textnormal{d}x\textnormal{d}{s}\;{\Big|}\;\mathcal{F}_{t}\right)+[pK+(p-1)]{\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}|y|^{p}\,\textnormal{d}x\textnormal{d}{s}\;{\Big|}\;\mathcal{F}_{t}\right)$ (55)
whence, by taking expectation and using Fubini’s theorem yield, we get
$\mathbb{E}\left(\left\lVert y(t)\right\rVert_{L^{p}(\mathcal{O})}^{p}\right)\leq\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))}^{p}+\left\lVert F\right\rVert_{L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))}^{p}+[pK+(p-1)]\int_{t}^{T}\mathbb{E}\left(\left\lVert y(s)\right\rVert_{L^{p}(\mathcal{O})}^{p}\right)\textnormal{d}{s},$
for all $t\in[0,T]$.
To finish, we apply backward Grönwall estimate (see Lemma˜A.1) to deduce
${\mathbb{E}}\left(\left\lVert y(t)\right\rVert_{L^{p}(\mathcal{O})}^{p}\right)\leq\exp([pK+(p-1)](T-t))\left(\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))}^{p}+\left\lVert F\right\rVert_{L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))}^{p}\right),$
that is
$\left\lVert y(t)\right\rVert_{L^{p}(\Omega;L^{p}(\mathcal{O}))}^{p}\leq\exp([pK+(p-1)](T-t))\left(\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))}^{p}+\left\lVert F\right\rVert_{L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))}^{p}\right).$
We take the power $1/p$ and we get
$\left\lVert y(t)\right\rVert_{L^{p}(\Omega;L^{p}(\mathcal{O}))}$ $\leq\exp((K+(p-1)/p)(T-t))\left(\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))}^{p}+\left\lVert F\right\rVert_{L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))}^{p}\right)^{1/p}$
$\leq\exp([K+1](T-t))\left(\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))}+\left\lVert F\right\rVert_{L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))}\right),$
where we have used that (p1)/p<1(p-1)/p<1. Taking $L^{\infty}$-norm in time yields the desired result. ∎
We present the proof of our main result.
Proof of Theorem˜1.4.
From Proposition˜3.1, we have that the following estimate holds
$\left\lVert y(t)\right\rVert_{L^{p}(\Omega;L^{p}(\mathcal{O}))}\leq\exp(CT)\left(\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))}+\left\lVert F\right\rVert_{L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))}\right)$ (56)
for all $t\in[0,T]$, where $y$ can be found from $(y,Y)$ solution to (1) and $C>0$ is a constant independent of $p$.
By Hölder inequality, the right-hand side of the above expression is bounded as
$R.H.S.$ $\leq\exp(CT)\left(|\mathcal{O}|^{1/p}\left\lVert y_{T}\right\rVert_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}+|\mathcal{O}|^{1/p}|T|^{1/p}\left\lVert F\right\rVert_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\right)$
$\leq\exp(C^{\prime}(1+T))\left(\left\lVert y_{T}\right\rVert_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}+\left\lVert F\right\rVert_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\right)$ (57)
for a constant $C^{\prime}>0$ depending on $\mathcal{O}$ but uniform with respect to $T$ and $p$. Estimate (57) and Lemma˜A.2 allows us to pass to the limit as $p\to+\infty$ in (56) and obtain
$\left\lVert y(t)\right\rVert_{L^{\infty}(\Omega;L^{\infty}(\mathcal{O}))}\leq\exp(CT)\left(\left\lVert y_{T}\right\rVert_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}+\left\lVert F\right\rVert_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\right)$
for every $t\in[0,T]$. Therefore,
$\left\lVert y\right\rVert_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\Omega;L^{\infty}(\mathcal{O})))}\leq\exp(CT)\left(\left\lVert y_{T}\right\rVert_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}+\left\lVert F\right\rVert_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\right).$
The result (6) then follows by interchanging the $L^{\infty}$-norms. ∎
4 Local well-posedness of semilinear equations
In this section, we present the proof of Theorem˜1.7. Hereinafter, we set $Q_{T}:=(0,T)\times\mathcal{O}$ and $\Sigma_{T}:=(0,T)\times\partial\mathcal{O}$ to abridge the notation.
We begin by stating the following result.
Proposition 4.1.
Let $p\in[2,+\infty)$, $(y,Y)$ be a weak solution to (1) and assume that $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))$ and $F\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))$. Then, $(y,Y)\in\mathscr{Y}_{p}$ and satisfies
$\|(y,Y)\|_{\mathscr{Y}_{p}}\leq C\left(\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}+\|F\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\right),$ (58)
for a constant $C>0$ only depending on $\mathcal{O}$, $T$, and $p$.
Proof.
The proof is a direct consequence of Hölder inequality, Theorem˜1.2, point (ii) of Theorem˜1.3, and Proposition˜2.6. ∎
Proof of Theorem˜1.7.
We will see that if the terminal data $y_{T}$ is small enough, then there is a unique solution $(y,Y)$ to (7) in $\mathscr{Y}_{p}$. We split the proof in two parts.
-Existence. By Taylor’s formula at second order, we have that
$\forall s\in\mathbb{R},\quad f(s)=f^{\prime}(0)s+s^{2}G(s),\quad$ (59)
where $G(s)=\int_{0}^{1}(1-\sigma)f^{\prime\prime}(\sigma s)\textnormal{d}{\sigma}$. For any $r\in(0,1)$, let
$\mathscr{Y}^{r}_{p}:=\{(y,Y)\in\mathscr{Y}_{p}:\|(y,Y)\|_{\mathscr{Y}_{p}}\leq r\},$ (60)
and consider the linear equation
$\begin{cases}\textnormal{d}{y}=-\left(\Delta y+\alpha y+\beta Y+f^{\prime}(0)y+\overline{y}^{2}G(\overline{y})\right)\textnormal{d}{t}+Y\textnormal{d}{W}(t)&\text{in }Q_{T},\\ y=0&\text{on }\Sigma_{T},\\ y(T)=y_{T}&\text{in }\mathcal{O},\end{cases}$ (61)
where $\overline{y}$ is the first entry of any given pair $(\overline{y},\overline{Y})\in\mathscr{Y}_{p}^{r}$ and $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))$. Note that this is a linearized version of (7).
Let us denote by $(y,Y)$ the solution to (61). We will check that the mapping
$\begin{split}\mathcal{N}:\mathscr{Y}_{p}^{r}&\longrightarrow\mathscr{Y}_{p}^{r}\\ (\overline{y},\overline{Y})&\longmapsto(y,Y)\end{split}$
is well-defined if we consider terminal data $y_{T}$ small enough. Since $f\in C^{\infty}(\mathbb{R})$, from point i) of Lemma˜A.3, we have that $|G(s)|\leq M$ for all $s\in[-2,2]$ where the constant $M>0$ only depends on $f$. Hence, defining $\overline{F}:=\overline{y}^{2}G(\overline{y})$ we have
$\|\overline{F}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$ $=\|\overline{y}^{2}G(\overline{y})\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$
$\leq M\|\overline{y}^{2}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}=M\|\overline{y}\|^{2}_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$
$\leq M\|(\overline{y},\overline{Y})\|_{\mathscr{Y}_{p}}^{2}\leq Mr^{2}<+\infty,$ (62)
for any $(\overline{y},\overline{Y})\in\mathscr{Y}_{p}^{r}$.
Thus, using Proposition˜4.1 with this particular $\overline{F}$ and estimate (62) yield that there is a solution $(y,Y)\in\mathscr{Y}_{p}$ to (61) such that
$\|(y,Y)\|_{\mathscr{Y}_{p}}$ $\leq C\left(\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}+Mr^{2}\right)$ (63)
for some $C>0$ only depending on $\mathcal{O}$, $T$ and $p$. Let us take
$r\leq\frac{1}{8CM},$
where $C$ is the constant appearing in (63) and fix $\delta>0$ such that $\delta\leq r/2C$. Thus for any $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))$ verifying $\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\leq\delta$ we have
$\|(y,Y)\|_{\mathscr{Y}_{p}}\leq\frac{r}{2}+\frac{r}{8}=\frac{5r}{8}\leq r.$ (64)
This shows that the map $\mathcal{N}:\mathscr{Y}_{p}^{r}\to\mathscr{Y}_{p}^{r}$ is well-defined.
To conclude, we will verify that $\mathcal{N}$ is a contraction on $\mathscr{Y}_{p}^{r}$ and use Banach fixed point theorem. By point ii) of Lemma˜A.3 with $\mathcal{I}=[-2,2]$, we have that for any $\overline{y}_{1}$, $\overline{y}_{2}$ taken from the pairs $(\overline{y}_{i},\overline{Y}_{i})\in\mathscr{Y}_{p}^{r}$, $i=1,2$,
$\|\overline{y}_{1}G(\overline{y}_{1})-\overline{y}_{2}G(\overline{y}_{2})\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$
$\qquad\leq M\|\overline{y}_{1}-\overline{y}_{2}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\left(\|\overline{y}_{1}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}+\|\overline{y}_{2}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\right)$
$\quad\qquad+M_{1}\|y_{2}\|^{2}_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\|\overline{y}_{1}-\overline{y}_{2}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))},$ (65)
for some positive constants $M,M_{1}$ only depending on $f$.
Denote by $(y_{i},Y_{i})$ the solution of (61) for the corresponding $\overline{y}_{i}$, $i=1,2$. From the linearity of (61), Proposition˜4.1 with $\overline{F}=\overline{y}_{1}G(\overline{y}_{1})-\overline{y}_{2}G(\overline{y}_{2})$ and $y_{T}\equiv 0$, and estimate (65), we can compute
$\|\mathcal{N}(\overline{y}_{1},\overline{Y}_{1})-\mathcal{N}(\overline{y}_{2},\overline{Y}_{2})\|_{\mathscr{Y}_{p}}=\|\mathcal{(}{y}_{1},{Y}_{1})-({y}_{2},{Y}_{2})\|_{\mathscr{Y}_{p}}$
$\qquad\leq CM\|\overline{y}_{1}-\overline{y}_{2}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\left(\|\overline{y}_{1}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}+\|\overline{y}_{2}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\right)$
$\quad\qquad+CM_{1}\|y_{2}\|^{2}_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\|\overline{y}_{1}-\overline{y}_{2}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$
$\qquad\leq 2CMr\|\overline{y}_{1}-\overline{y}_{2}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}+CM_{1}r^{2}\|\overline{y}_{1}-\overline{y}_{2}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}.$ (66)
From (66) and decreasing (if necessary) the value of $r$ and taking $r\leq\min\{\frac{1}{8CM},\frac{1}{\sqrt{4CM_{1}}}\}$, we get
$\|\mathcal{N}(\overline{y}_{1},\overline{Y}_{1})-\mathcal{N}(\overline{y}_{2},\overline{Y}_{2})\|_{\mathscr{Y}_{p}}\leq\frac{1}{2}\|\overline{y}_{1}-\overline{y}_{2}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\leq\frac{1}{2}\|(\overline{y}_{1},\overline{Y}_{1})-(\overline{y}_{2},\overline{Y}_{2})\|_{\mathscr{Y}_{p}}.$
Thus $\mathcal{N}$ is a contraction on the set $\mathscr{Y}_{p}^{r}$. A direct application of Banach fixed point theorem yields that there is a unique fixed point for $\mathcal{N}$ which is a weak solution to (7) thanks to (59) and that belongs to $\mathscr{Y}_{p}^{r}$ since (64) holds.
-Uniqueness. Let $\delta>0$ be fixed and small enough as in the previous part. Let $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))$ be such that $\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\leq\delta$ and assume that there are two weak solutions $(y_{1},Y_{1})$ and $(y_{2},Y_{2})$ to (7) for this same data that belong to some ball $\mathscr{Y}_{p}^{r}$ (recall (60)). Defining $(y,Y):=(y_{1}-y_{2},Y_{1}-Y_{2})$, it is not difficult to see that these new variables verify
$\begin{cases}\textnormal{d}{y}=-\left(\Delta y+\alpha y+\beta Y+\left(f(y_{1})-f(y_{2})\right)\right)\textnormal{d}{t}+Y\textnormal{d}{W}(t)&\text{in }Q_{T},\\ y=0&\text{on }\Sigma_{T},\\ y(T)=0&\text{in }\mathcal{O}.\end{cases}$
Arguing as we did for obtaining (55) and taking expectation, we can deduce
$\mathbb{E}\left(\left\lVert y(t)\right\rVert_{L^{p}(\mathcal{O})}^{p}\right)\leq{\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}|f(y_{1})-f(y_{2})|^{p}\,\textnormal{d}x\textnormal{d}{s}\right)+C{\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}|y|^{p}\,\textnormal{d}x\textnormal{d}{s}\right),$
where $C>0$ is a constant only depending on $p$, $\alpha$ and $\beta$. Since $(y_{i},Y_{i})\in\mathscr{Y}_{p}^{r}$, $i=1,2$, and using the fact that $f$ is locally Lipschitz on [r,r][-r,r], there is a constant $C^{\prime}>0$ depending on $r$ such that $|f(s_{1})-f(s_{2})|\leq C^{\prime}|s_{1}-s_{2}|$, $s_{1},s_{2}\in\mathbb{[}-r,r]$, therefore
$\mathbb{E}\left(\left\lVert y(t)\right\rVert_{L^{p}(\mathcal{O})}^{p}\right)\leq C^{\prime\prime}{\mathbb{E}}\left(\int_{t}^{T}\int_{\mathcal{O}}|y|^{p}\,\textnormal{d}x\textnormal{d}{s}\right),$
for a constant $C^{\prime\prime}>0$. A standard Grönwall argument using Lemma˜A.1, yields that $y_{1}=y_{2}$ a.e $(t,x)\in Q_{T}$, a.s. This ends the proof. ∎
5 Applications to controllability
The goal of this section is to present some applications to the control of backward stochastic parabolic equations, first in the linear case and then in the nonlinear framework. In particular, we will prove Proposition˜1.9 and Theorem˜1.11.
5.1 The linear case
We consider the following backward stochastic heat equation
$\begin{cases}\textnormal{d}{y}=-\left(\Delta y+\alpha y+\chi_{\mathcal{O}_{0}}h\right)\textnormal{d}{t}+Y\textnormal{d}{W}(t)&\text{in }Q_{T},\\ y=0&\text{on }\Sigma_{T},\\ y(T)=y_{T}&\text{in }\mathcal{O}.\end{cases}$ (67)
In (67), we recall that $(y,Y)$ the state variable while $h$ is the control variable located in $\mathcal{O}_{0}$ where $\mathcal{O}_{0}$ is a nonempty open subset of $\mathcal{O}$.
We now present the proof of the null-controllability result of Proposition˜1.9.
Proof of Proposition˜1.9.
We take $p\in(2,+\infty)$ so $p^{\prime}\in(1,2)$.
The adjoint of (67) is given by the random heat equation
$\begin{cases}\textnormal{d}{q}=(\Delta q+\alpha q)\textnormal{d}t&\text{in }Q_{T},\\ q=0&\text{on }\Sigma_{T},\\ q(0)=q_{0}&\text{in }\mathcal{O}.\end{cases}$ (68)
Note that the randomness in equation (68) appears only through the adapted coefficient $\alpha\in L_{\mathbb{F}}^{\infty}(0,T;L^{\infty}(\mathcal{O}))$ and the $\mathcal{F}_{0}$-measurable initial datum $q_{0}\in L^{p^{\prime}}_{\mathcal{F}_{0}}(\Omega;L^{p^{\prime}}(\mathcal{O}))$. Moreover, for such coefficients, the weak solution $t\mapsto q(t,\cdot)\in L^{2}(\mathcal{O})$ to (68) is $\mathcal{F}_{t}$-measurable for every $t\in[0,T]$.
For $\omega\in\Omega$ fixed, we have from [FZ00, Proposition 3.2]
$\left\lVert q(\omega,T)\right\rVert_{L^{2}(\mathcal{O})}\leq C\left\lVert q(\omega)\right\rVert_{L^{1}((0,T)\times\mathcal{O}_{0})},$ (69)
from which we directly deduce, by using $L^{2}(\mathcal{O})\hookrightarrow L^{p^{\prime}}(\mathcal{O})$ and $L^{p^{\prime}}((0,T)\times\mathcal{O}_{0}))\hookrightarrow L^{1}((0,T)\times\mathcal{O}_{0}))$,
$\left\lVert q(\omega,T)\right\rVert_{L^{p^{\prime}}(\mathcal{O})}\leq C\left\lVert q(\omega)\right\rVert_{L^{p^{\prime}}((0,T)\times\mathcal{O}_{0})}.$
By taking the power $p^{\prime}$ and by taking the expectation we finally get
$\mathbb{E}\Big[\left\lVert q(T)\right\rVert_{L^{p^{\prime}}(\mathcal{O})}^{p^{\prime}}\Big]\leq C\mathbb{E}\Big[\left\lVert q\right\rVert_{L^{p^{\prime}}((0,T)\times\mathcal{O}_{0})}^{p^{\prime}}\Big].$
This rewrites as
$\left\lVert q(T)\right\rVert_{L^{p^{\prime}}_{\mathcal{F}_{T}}(\Omega;L^{p^{\prime}}(\mathcal{O}))}\leq C\left\lVert q\right\rVert_{L^{p^{\prime}}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}_{0}))}.$ (70)
Note that the constant $C>0$ in (70) is of the form $\exp(C/T)$ for $T\in(0,1)$ for some $C$ depending on $\mathcal{O}$, $\mathcal{O}_{0}$, $\alpha$ and $p$ because the observability constant $C$ in (69) appearing in the deterministic observability inequality is of the form $\exp(C/T)$.
Now, we will follow an argument given by [TZ09], adapted to the $L^{p^{\prime}}$-setting. For any $y_{T}\in L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))$, the goal is to find a control $h\in L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}_{0}))$ such that the solution $y$ of (67) satisfies $y(0)=0$ a.s.
We introduce the following linear subspace of $L_{\mathbb{F}}^{p^{\prime}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0}))$,
[…]
Then we define the following linear functional on $\mathcal{X}$ by
$\mathcal{L}(q)=-\mathbb{E}\int_{\mathcal{O}}y_{T}q(T)\,\textnormal{d}x.$
The observability estimate (70) leads to the fact that $\mathcal{L}$ is a bounded linear functional on $\mathcal{X}$ because
$\left|\mathbb{E}\int_{\mathcal{O}}y_{T}q(T)\,\textnormal{d}x\right|\leq\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))}\left\lVert q(T)\right\rVert_{L^{p^{\prime}}_{\mathcal{F}_{T}}(\Omega;L^{p^{\prime}}(\mathcal{O}))}\leq C\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))}\left\lVert q\right\rVert_{L^{p^{\prime}}_{\mathbb{F}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0}))}.$ (71)
By the Hahn-Banach theorem, $\mathcal{L}$ can be extended to a bounded linear functional on $L_{\mathbb{F}}^{p^{\prime}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0}))$ with
$\|\mathcal{L}\|_{\mathcal{L}(L_{\mathbb{F}}^{p^{\prime}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0})),{\mathbb{R}})}\leq C\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))},$ (72)
where $C$ is the same as in (70). Now, we use the following fact coming from [LYZ12, Corollary 2.3]
$(L_{\mathbb{F}}^{p^{\prime}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0}))^{\prime}=L_{\mathbb{F}}^{p}(0,T;L^{p}(\mathcal{O}_{0})).$
This means that there exists $h\in L_{\mathbb{F}}^{p}(0,T;L^{p}(\mathcal{O}_{0}))$ such that
$\mathcal{L}(q)=\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hq\,\textnormal{d}x\textnormal{d}t\qquad\forall q\in L^{p^{\prime}}_{\mathbb{F}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0})),$ (73)
in particular for $q$ solution to (68)
$\mathcal{L}(q)=-\mathbb{E}\int_{\mathcal{O}}y_{T}q(T)\,\textnormal{d}x=\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hq\,\textnormal{d}x\textnormal{d}t\qquad\forall q_{0}\in L^{p^{\prime}}_{\mathcal{F}_{0}}(\Omega;L^{p^{\prime}}(\mathcal{O})).$ (74)
Now we use Itô’s formula to get
$\textnormal{d}(yq)=y\textnormal{d}{q}+q\textnormal{d}{y}+\textnormal{d}{y}\textnormal{d}{q}.$
So we get
$\mathbb{E}\int_{\mathcal{O}}y_{T}q(T)\,\textnormal{d}x-\mathbb{E}\int_{\mathcal{O}}y(0)q_{0}\,\textnormal{d}x=\mathbb{E}\int_{Q_{T}}y\textnormal{d}{q}+q\textnormal{d}{y}+\textnormal{d}{y}\textnormal{d}{q}=\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hq\,\textnormal{d}x\textnormal{d}t.$
Therefore, we have
$\mathbb{E}\int_{\mathcal{O}}y(0)q_{0}\,\textnormal{d}x=0\qquad\forall q_{0}\in L^{p^{\prime}}_{\mathcal{F}_{0}}(\Omega;L^{p^{\prime}}(\mathcal{O})).$
This implies that $y(0)=0$ in $\mathcal{O}$ a.s. Moreover, we have from (72) and (73) that
$\|h\|_{L_{\mathbb{F}}^{p}(0,T;L^{p}(\mathcal{O}_{0}))}\leq C\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))},$
where
$C=C(T)\leq\exp(C/T),$
hence the expected bound on $h$ in (10). This concludes the proof in the case $p\in(2,+\infty)$.
Now we deal with the more subtle case $p=+\infty$. The difficulty is that one cannot use [LYZ12, Corollary 2.3] because the so-called Radon-Nikodym property is not satisfied by $L^{\infty}$, in particular $(L_{\mathbb{F}}^{1}(0,T;L^{1}(\mathcal{O}_{0}))^{\prime}$ is not equal to $L_{\mathbb{F}}^{\infty}(0,T;L^{\infty}(\mathcal{O}_{0}))$. We start from the $L^{1}$-observability estimate for solution $q$ to (68)
$\left\lVert q(T)\right\rVert_{L^{1}_{\mathbb{F}}(\Omega;L^{1}(\mathcal{O}))}\leq C\left\lVert q\right\rVert_{L^{1}_{\mathbb{F}}(0,T;L^{1}(\mathcal{O}_{0}))},$ (75)
that is an easy consequence of (69). We then introduce the following linear subspace of $L_{\mathcal{F}}^{1}(0,T;L^{1}(\mathcal{O}_{0}))$,
[…]
Then we define the following linear functional on $\mathcal{X}$ by
$\mathcal{L}(q)=-\mathbb{E}\int_{\mathcal{O}}y_{T}q(T)dx.$
The observability estimate (75) leads to the fact that $\mathcal{L}$ is a bounded linear functional on $\mathcal{X}$. By the Hahn-Banach theorem, $\mathcal{L}$ can be extended to a bounded linear functional on $L_{\mathcal{F}}^{1}(0,T;L^{1}(\mathcal{O}_{0}))$ with
$\|\mathcal{L}\|_{\mathcal{L}(L_{\mathbb{F}}^{1}(0,T;L^{1}(\mathcal{O}_{0})),{\mathbb{R}})}\leq C\left\lVert y_{T}\right\rVert_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))},$ (76)
where $C$ is the same as in (75). In particular, $\mathcal{L}$ is a bounded linear functional on $L_{\mathbb{F}}^{2}(0,T;L^{2}(\mathcal{O}_{0}))$ because $L_{\mathbb{F}}^{2}(0,T;L^{2}(\mathcal{O}_{0}))\hookrightarrow L_{\mathbb{F}}^{1}(0,T;L^{1}(\mathcal{O}_{0}))$, therefore we can use the following fact coming from [LYZ12, Corollary 2.3],
$(L_{\mathbb{F}}^{2}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O}_{0})))^{\prime}=L_{\mathbb{F}}^{2}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O}_{0}))).$
This means that there exists $h\in L_{\mathbb{F}}^{2}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O}_{0})))$ such that
$\mathcal{L}(q)=\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hqdxdt\qquad\forall q\in L^{2}_{\mathbb{F}}(0,T;L^{2}(\mathcal{O}_{0})),$ (77)
in particular for $q$ solution to (68)
$\mathcal{L}(q)=-\mathbb{E}\int_{\mathcal{O}}y_{T}q(T)dx=\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hqdxdt\qquad\forall q_{0}\in L^{2}_{\mathcal{F}_{0}}(\Omega;L^{2}(\mathcal{O})).$
As before we easily deduce by Itô’s formula that $y(0)=0$ in $\mathcal{O}$ a.s.
Our goal is then to prove that $h\in L_{\mathbb{F}}^{\infty}(0,T;L^{\infty}(\mathcal{O}_{0}))$ and
$\|h\|_{L_{\mathbb{F}}^{\infty}(0,T;L^{\infty}(\mathcal{O}_{0}))}\leq 2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}.$ (78)
From (76) and (77), we deduce that
$\left|\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hqdxdt\right|\leq C\|q\|_{L^{1}_{\mathbb{F}}(0,T;L^{1}(\mathcal{O}))}\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\quad\forall q\in L^{2}_{\mathbb{F}}(0,T;L^{2}(\mathcal{O}_{0})).$ (79)
Then we take
$q=\text{sign}(h)\mathbf{1}_{\{|h|>2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\}}\mathbf{1}_{\mathcal{O}_{0}},$
so the left hand side of (79) is bounded from below by
$2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\mu_{\Omega,[0,T],\mathcal{O}_{0}}(|h|>2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}),$
while the right hand side of (79) translates into
$C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\mu_{\Omega,[0,T],\mathcal{O}_{0}}(|h|>2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}),$
where $\mu_{\Omega,[0,T],\mathcal{O}_{0}}$ denotes the product measure on $\Omega\times[0,T]\times\mathcal{O}$, so we get
$2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\mu_{\Omega,[0,T],\mathcal{O}_{0}}(|h|>2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))})\\ \leq C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\mu_{\Omega,[0,T],\mathcal{O}_{0}}(|h|>2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}).$
That is a contradiction unless we have
$\mu_{\Omega,[0,T],\mathcal{O}_{0}}(|h|>2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))})=0.$
This exactly means that we have
$|h|\leq 2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\quad\text{a.e. }(t,x)\in Q_{T},\ \text{a.s.}$
This concludes the proof of (78), then the one of Proposition˜1.9. ∎
5.2 The nonlinear control problem
The proof of Theorem˜1.11 relies on the so-called source term method, originally developed in [LTT13] for the deterministic setting and later extended to SPDEs in [HLP22]. While [HLP22] already discusses this method for BSPDEs, here we provide a more comprehensive treatment and extend the analysis to our setting.
Let us consider a slightly more general version of the linear equation (67), namely
$\begin{cases}\textnormal{d}{y}=-\left(\Delta y+\alpha y+\chi_{\mathcal{O}_{0}}h+S\right)\textnormal{d}{t}+Y\textnormal{d}{W}(t)&\text{in }Q_{T},\\ y=0&\text{on }\Sigma_{T},\\ y(T)=y_{T}&\text{in }\mathcal{O},\end{cases}$ (80)
where $S$ is an external source term that will be taken from a suitable functional space and $h$ is a control.
From Proposition˜4.1, we can readily deduce that if the control $h\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}_{0}))$ and the source term $S\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))$, then (80) is well-posed and $(y,Y)\in\mathscr{Y}_{p}$ for any $p\in[2,+\infty)$ (recall (8)). On the other hand, Proposition˜1.9 already tells us that if $S\equiv 0$, then (80) is null-controllable in $L^{\infty}$. Our first aim is to see that if $S$ belongs to an appropriate space, then (80) remains null-controllable in $L^{\infty}$.
To this end, we fix $M\geq 1$ such that
$e^{C/T}\leq Me^{M/T},$ (81)
where $C$ is the constant provided by Proposition˜1.9 (with $p=+\infty$). For some parameters $Q\in(1,\sqrt{2})$ and $R>Q^{2}/(2-Q^{2})$, we define
$\rho(t)$ $:=\exp\left(-\frac{(1+R)Q^{2}M}{(Q-1)t}\right),$ (82)
$\rho_{0}(t)$ $:=\exp\left(-\frac{RM}{(Q-1)t}\right),$ (83)
which are strictly increasing and vanish as $t\to 0^{+}$.
Using these functions, we introduce the following weighted spaces
$\mathscr{S}$ $:=\left\{S\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O})):\ \frac{S}{\rho}\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))\right\},$ (84)
$\mathscr{E}$ $:=\left\{y\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O})):\ \frac{y}{\rho_{0}}\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))\right\},$ (85)
$\mathscr{H}$ $:=\left\{h\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}_{0})):\ \frac{h}{\rho_{0}}\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}_{0}))\right\}.$ (86)
We have the following result.
Proposition 5.1.
For every $S\in\mathscr{S}$ and $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\Omega))$, there is a control $h\in\mathscr{H}$ such that $y$ the first component of the solution $(y,Y)\in\mathscr{Y}_{p}$ to (80) satisfies $y\in\mathscr{E}$. Moreover, the exists a constant $C>0$ independent of $S$ and $y_{T}$ such that
$\|y/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}+\|h/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}+\|Y\|_{L^{p}_{\mathbb{F}}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O})))}$
$\qquad\leq Ce^{C/T}\left(\left\lVert y_{T}\right\rVert_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}+\left\lVert S/\rho\right\rVert_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\right),$ (87)
Proof.
The proof follows the same steps of [HLP22, Proposition 2.5] with some adaptations to the backward case and the $L^{\infty}$-setting (cf. [LE 20]).
Let $T_{k}:=TQ^{-k}$ where we recall that $Q\in(1,\sqrt{2})$ is a given parameter. Let $a_{0}:=y_{T}$ and for $k\in\mathbb{N}$ we define $a_{k+1}:=y_{S}(T_{k+1},\cdot)$ where $y_{S}$ is the first component of $(y_{S},Y_{S})$ the solution to the uncontrolled system
$\begin{cases}\textnormal{d}{y_{S}}=-\left(\Delta y_{S}+\alpha y_{S}+S\right)\textnormal{d}{t}+Y_{S}\textnormal{d}{W}(t)&\text{in }(T_{k+1},T_{k})\times\mathcal{O},\\ y_{S}=0&\text{on }(T_{k+1},T_{k})\times\mathcal{O},\\ y_{S}(T_{k})=0&\text{in }\mathcal{O},\end{cases}$ (88)
We observe that by Proposition˜4.1, $y_{S}$ is a continuous adapted function in $T_{k+1},T_{k}]$, $a_{k+1}$ is $\mathcal{F}_{T_{k+1}}$ measurable and belongs to $L^{\infty}(\Omega\times\mathcal{O})$. In particular, by estimate (58)
$\|a_{k+1}\|_{L^{\infty}(\Omega\times\mathcal{O})}\leq\|y\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}))}\leq\|(y,Y)\|_{\mathscr{Y}_{p}}\leq C\|S\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}))}.$ (89)
Let $\{a_{k}\}_{k\in\mathbb{N}}$ be the sequence defined by the above procedure. For each $k\in\mathbb{N}$, consider the controlled equation (without a source term) given by
$\begin{cases}\textnormal{d}{y_{H}}=-\left(\Delta y_{H}+\alpha y_{H}+\chi_{\mathcal{O}_{0}}h_{k}\right)\textnormal{d}{t}+Y_{H}\textnormal{d}{W}(t)&\text{in }(T_{k+1},T_{k})\times\mathcal{O},\\ y_{H}=0&\text{on }(T_{k+1},T_{k})\times\partial\mathcal{O},\\ y_{H}(T_{k})=a_{k}&\text{in }\mathcal{O}.\end{cases}$ (90)
This system is well-posed (in the sense of Proposition˜4.1) for each $k\in\mathbb{N}$ due to the properties of $a_{k}$ mentioned above. In view of Proposition˜1.9, we can obtain controls $h_{k}\in L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}_{0}))$ such that $y_{H}(T_{k+1})=0$. Moreover, using (81) and the precise estimate (10), it can be deduced that
$\|h_{k}\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}_{0}))}\leq Me^{\frac{M}{T_{k}-T_{k+1}}}\|a_{k}\|_{L^{\infty}_{\mathcal{F}_{T_{k}}}(\Omega;L^{\infty}(\mathcal{O}))},\qquad\forall k\in\mathbb{N}.$ (91)
In particular, for $k=0$, we have
$\|h_{0}\|_{L^{\infty}_{\mathbb{F}}(T_{1},T;L^{\infty}(\mathcal{O}_{0}))}\leq Me^{\frac{MQ}{T(Q-1)}}\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))},$ (92)
and, from the definition of $\rho_{0}$ (see (83)), we see that $\rho_{0}(t_{1})<\rho_{0}(t_{2})$ for every $0\leq t_{1}<t_{2}\leq T$, whence
$\|h_{0}/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(T_{1},T;L^{\infty}(\mathcal{O}_{0}))}\leq\rho_{0}^{-1}(T_{1})Me^{\frac{MQ}{T(Q-1)}}\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}$ (93)
On the other hand, putting together (89) and (91) yields
$\|h_{k+1}\|_{L^{\infty}_{\mathbb{F}}(T_{k+2},T_{k+1};L^{\infty}(\mathcal{O}))}\leq CMe^{\frac{M}{T_{k+1}-T_{k+2}}}\|S\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}))}$ (94)
Similar as above, from (82) we have $\rho(t_{1})<\rho(t_{2})$ for every $0\leq t_{1}<t_{2}\leq T$ and therefore
$\|h_{k+1}\|_{L^{\infty}_{\mathbb{F}}(T_{k+2},T_{k+1};L^{\infty}(\mathcal{O}))}\leq CMe^{\frac{M}{T_{k+1}-T_{k+2}}}\rho(T_{k})\|S/\rho\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}))}.$ (95)
We note that by definitions (82) and (83), the identity $e^{\frac{M}{T_{k+1}-T_{k+2}}}\rho(T_{k})=\rho_{0}(T_{k+2})$ holds. Thus,
$\|h_{k+1}\|_{L^{\infty}_{\mathbb{F}}(T_{k+2},T_{k+1};L^{\infty}(\mathcal{O}))}\leq CM\rho_{0}(T_{k+2})\|S/\rho\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}))},$ (96)
and since $\rho_{0}(T_{k+2})\leq\rho(t)$ for all $t\in(T_{k+2},T_{k+1})$
$\|h_{k+1}/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(T_{k+2},T_{k+1};L^{\infty}(\mathcal{O}))}\leq CM\|S/\rho\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}))},\qquad\forall k\in\mathbb{N}.$ (97)
Let us define $h(t,\cdot):=\sum_{k\in\mathbb{N}}1_{\left(T_{k+1},T_{k}\right)}(t)h_{k}(t,\cdot)$, where each control $h_{k}$ satisfies (97) for $k\geq 1$ and (93) if $k=0$. Defined in this way, since each interval $(T_{k+1},T_{k})$ is disjoint and $h_{k}$ are independent processes for each $k\in\mathbb{N}$, we have that $h\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))$. Moreover, by (93), (97), and using the fact that the essential supremum over disjoint intervals equals the supremum of the individual norms, we get
$\|h/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$ $=\sup_{k\in\mathbb{N}}\|h_{k}/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}))}$
$\leq\rho_{0}^{-1}(T_{1})Me^{\frac{MQ}{T(Q-1)}}\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}+CM\sup_{k\in\mathbb{N}}\|S/\rho\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}))}$
$=C^{\prime}e^{C^{\prime}/T}\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}+C^{\prime}\|S/\rho\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$ (98)
for some constant $C^{\prime}>0$ uniform with respect to $y_{T}$ and $S$.
In a similar way, we can build the solution to (80) by gluing together $y_{S}$ and $y_{H}$, the solutions to (88) and (90), respectively. Indeed, from Proposition˜4.1 and recalling (8), we have for each time interval $(T_{k+1},T_{k})$
$\|y_{S}\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}))}\leq C\|S\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}))}$ (99)
and
$\|y_{H}\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}))}\leq C\left(\|a_{k}\|_{L^{\infty}_{\mathcal{F}_{T_{k}}}(\Omega;L^{\infty}(\mathcal{O}))}+\|h_{k}\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}))}\right).$ (100)
We note that, by construction, $y(t,\cdot)=y_{S}(t,\cdot)+y_{H}(t,\cdot)$ is a continuous adapted function in each subinterval $(T_{k+1},T_{k})$ and, moreover, it is continuous at the junction times $T_{k}$, $k\in\mathbb{N}$. Furthermore, we can deduce from (93), (97), (99), and (100), using an argument similar to that for the control, that
$\|y/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$ $\leq\|y_{H}/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}+\|y_{S}/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$
$=\sup_{k\in\mathbb{N}}\left(\|y_{H}/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}))}+\|y_{S}/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}))}\right)$
$\leq C\rho_{0}^{-1}(T_{1})Me^{\frac{MQ}{T(Q-1)}}\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}+CM^{-1}\sup_{k\in\mathbb{N}}\|S/\rho\|_{L^{\infty}_{\mathbb{F}}(T_{k+1},T_{k};L^{\infty}(\mathcal{O}))}$
$=C^{\prime}e^{C^{\prime}/T}\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}+C^{\prime\prime}\|S/\rho\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}.$ (101)
for some constant $C^{\prime\prime}>0$ independent of $S$.
Estimate (87) follows from (98), (101), and by choosing an appropriate constant $C>0$, which is uniform with respect to $S$ and $y_{T}$. To add the estimate on $Y$ in (87) we use (58) and the fact that $\rho$ is bounded in $[0,T]$. This ends the proof. ∎
Remark 5.2.
If the component $y$ of the solution $(y,Y)\in\mathscr{Y}_{p}$ of (80) satisfies $y\in\mathscr{E}$, then $y(0,\cdot)=0$ in $\Omega$ a.s.
We are now in position to prove our main controllability result.
Proof of Theorem˜1.11.
Let us define the following Banach space
$\mathscr{Z}_{p}=\mathscr{E}\times L^{p}_{\mathbb{F}}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O}))).$
By Taylor’s formula at second order, we have that
$\forall s\in\mathbb{R},\quad f(s)=f^{\prime}(0)s+s^{2}G(s),\quad$ (102)
where $G(s)=\int_{0}^{1}(1-\sigma)f^{\prime\prime}(\sigma s)\textnormal{d}{\sigma}$. For any $r\in(0,1)$, let
$\mathscr{Z}^{r}_{p}:=\{(y,Y)\in\mathscr{Z}_{p}:\|y\|_{\mathscr{E}}+\|Y\|_{L^{p}_{\mathbb{F}}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O})))}\leq r\},$ (103)
and consider the linear controlled equation
$\begin{cases}\textnormal{d}{y}=-\left(\Delta y+\alpha y+f^{\prime}(0)y+\overline{y}^{2}G(\overline{y})+\chi_{\mathcal{O}_{0}}h\right)\textnormal{d}{t}+Y\textnormal{d}{W}(t)&\text{in }Q_{T},\\ y=0&\text{on }\Sigma_{T},\\ y(T)=y_{T}&\text{in }\mathcal{O},\end{cases}$ (104)
where $\overline{y}$ is the first entry of any given pair $(\overline{y},\overline{Y})\in\mathscr{Z}_{p}^{r}$ and $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))$.
Let us define the following multi-valued contraction mapping
$\mathcal{N}:(\overline{y},\overline{Y})\in\mathscr{Z}_{p}^{r}\to 2^{\mathscr{Z}_{p}^{r}},$ (105)
such that for every $(\overline{y},\overline{Y})\in\mathscr{Z}_{p}^{r}$, $(y,Y)$ belongs to $\mathcal{N}((\overline{y},\overline{Y}))$, if there exists $h\in\mathscr{H}$ such that $(y,Y)\in\mathscr{Z}_{p}^{r}$ is the solution of (104) and estimate (87) holds.
Since $f\in C^{\infty}(\mathbb{R})$, from point i) of Lemma˜A.3, we have that $|G(s)|\leq M$ for all $s\in[-2,2]$ where the constant $M>0$ only depends on $f$. Moreover, we have from the definitions of the weights $\rho$ and $\rho_{0}$ that there exists $C>0$ such that $|(\rho_{0}^{2}/\rho)(t)|\leq C$ for every $t\in[0,T]$. Hence, defining $\overline{S}:=\overline{y}^{2}G(\overline{y})$ we have
$\|\overline{S}/\rho\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$ $=\|\overline{y}^{2}G(\overline{y})/\rho\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$
$\leq CM\|\overline{y}^{2}/\rho_{0}^{2}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$
$\leq CM\|\overline{y}/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}^{2}\leq CMr^{2}<+\infty,$ (106)
for any $(\overline{y},\overline{Y})\in\mathscr{Z}_{p}^{r}$.
Thus, using Proposition˜5.1 with this particular $\overline{S}$ and estimate (87) yield that there is a controlled trajectory solution $((y,Y),h)\in\mathscr{Z}_{p}\times\mathscr{H}$ to (104) such that
$\|y\|_{\mathscr{E}}+\|h\|_{\mathcal{H}}+\|Y\|_{L^{p}_{\mathbb{F}}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O})))}$ $\leq C\left(\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}+Mr^{2}\right)$ (107)
for some $C>0$ only depending on $\mathcal{O}$, $T$ and $p$. Let us take
$r\leq\frac{1}{8CM},$
where $C$ is the constant appearing in (107) and fix $\delta>0$ such that $\delta\leq r/2C$. Thus for any $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))$ verifying $\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\leq\delta$ we have
$\|y\|_{\mathscr{E}}+\|h\|_{\mathcal{H}}+\|Y\|_{L^{p}_{\mathbb{F}}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O})))}\leq\frac{r}{2}+\frac{r}{8}=\frac{5r}{8}\leq r.$ (108)
This shows that for every $(y,Y)\in\mathscr{Z}_{p}^{r}$, $\mathcal{N}(y,Y)$ is nonempty.
Moreover, for every $(y,Y)\in\mathscr{Z}_{p}^{r}$, from the estimate (87) it is straightforward to check that that $\mathcal{N}(y,Y)$ is a bounded and closed set of $\mathscr{Z}_{p}^{r}$.
To conclude, we will verify that $\mathcal{N}$ is a multi-valued contraction mapping on the Banach space $\mathscr{Z}_{p}^{r}$ and use the multi-valued Banach fixed point theorem from [NAD69, Theorem 5]. By point ii) of Lemma˜A.3 with $\mathcal{I}=[-2,2]$, we have that for any $\overline{y}_{1}$, $\overline{y}_{2}$ taken from the pairs $(\overline{y}_{i},\overline{Y}_{i})\in\mathscr{Z}_{p}^{r}$, $i=1,2$,
$\|(\overline{y}_{1}G(\overline{y}_{1})-\overline{y}_{2}G(\overline{y}_{2}))/\rho\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$
$\qquad\leq CM\|(\overline{y}_{1}-\overline{y}_{2})/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\left(\|\overline{y}_{1}/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}+\|\overline{y}_{2}/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\right)$
$\quad\qquad+CM_{1}\|y_{2}/\rho_{0}\|^{2}_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\|(\overline{y}_{1}-\overline{y}_{2})/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))},$ (109)
for some positive constants $M,M_{1}$ only depending on $f$. Above, we have used once again that $|(\rho_{0}^{2}/\rho)(t)|\leq C$ for every $t\in[0,T]$.
Denote by $(y_{i},Y_{i})\in\mathcal{N}(\overline{y_{i}},\overline{Y_{i}})$ the solution of (104) for the corresponding $\overline{y}_{i}$, $i=1,2$. From the linearity of (104), Proposition˜5.1 with $\overline{S}=\overline{y}_{1}G(\overline{y}_{1})-\overline{y}_{2}G(\overline{y}_{2})$ and $y_{T}\equiv 0$, and estimate (109), we can compute
$\|\mathcal{(}{y}_{1},{Y}_{1})-({y}_{2},{Y}_{2})\|_{\mathscr{Z}_{p}}$
$\qquad\leq CM\|(\overline{y}_{1}-\overline{y}_{2})/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\left(\|\overline{y}_{1}/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}+\|\overline{y}_{2}/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\right)$
$\quad\qquad+CM_{1}\|y_{2}/\rho_{0}\|^{2}_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\|(\overline{y}_{1}-\overline{y}_{2})/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}$
$\qquad\leq 2CMr\|(\overline{y}_{1}-\overline{y}_{2})/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}+CM_{1}r^{2}\|(\overline{y}_{1}-\overline{y}_{2})/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}.$ (110)
From (66) and decreasing (if necessary) the value of $r$ and taking $r\leq\min\{\frac{1}{8CM},\frac{1}{\sqrt{4CM_{1}}}\}$, we get
$\|\mathcal{(}{y}_{1},{Y}_{1})-({y}_{2},{Y}_{2})\|_{\mathscr{Z}_{p}}\leq\frac{1}{2}\|(\overline{y}_{1}-\overline{y}_{2})/\rho_{0}\|_{L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))}\leq\frac{1}{2}\|(\overline{y}_{1},\overline{Y}_{1})-(\overline{y}_{2},\overline{Y}_{2})\|_{\mathscr{Z}_{p}}.$
Thus $\mathcal{N}$ is a multi-valued contraction mapping on the Banach space $\mathscr{Z}_{p}^{r}$. Hence there is a fixed point for $\mathcal{N}$ which is a weak solution to (11) thanks to (102) and that belongs to $\mathscr{Z}_{p}^{r}$ since (108) holds. ∎
Remark 5.3.
The following comments are in order.
  • In the proof of Theorem˜1.11, we employed a multi-valued fixed point theorem, in contrast to the standard Banach fixed point argument used in Theorem˜1.7. The main reason for this is that Proposition˜5.1 guarantees only the existence of a control that steers the solution of (80) to zero at $t=0$. Similar strategies based on multi-valued fixed point theorems have been used in the control literature; see, for instance, the use of Kakutani’s fixed point theorem in [FZ00].
  • On the other hand, although the control produced in Theorem˜1.11 is not unique, the corresponding controlled trajectory of (11) is. Once a control steering the solution to zero is fixed, uniqueness follows by an argument similar to the one at the end of the proof of Theorem˜1.7.
6 On the optimality of the regularity
In this section, we provide additional remarks on the regularity results established in Theorems 1.3 and 1.4. By combining these results, we deduce that if $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))$ and $F\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))$, then the solution to (1) satisfies
$(y,Y)\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}))\times L^{p}_{\mathbb{F}}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O})))$ (111)
for every $p\geq 2$. However, from the proof of Theorem˜1.4, which itself relies on Proposition˜3.1, it is not evident whether the regularity of $Y$ (in $\Omega$) can be improved to $L^{\infty}$. In fact, the term involving $Y$ was explicitly removed in the derivation of the estimates (see, for example, inequality (55)). Therefore, the purpose of this section is to present an example showing that such an improvement is, in general, not possible in the simpler case of backward ordinary stochastic differential equations (BSDE).
To this end, let $T>0$ be a fixed deterministic time and consider the BSDE
$\begin{cases}\textnormal{d}y=\lambda y\textnormal{d}t+Y\textnormal{d}W(t)\quad t\in(0,T),\\ y(T)=y_{T},\end{cases}$ (112)
where $\lambda>0$ is a constant. Note that this equation can be interpreted as a projection of the linear stochastic system (1) onto a suitable subspace of $L^{2}$.
By standard well-posedness results for BSDEs (see, for instance, [YZ99, Chapter 7, Theorem 2.2]), equation (112) admits a unique adapted solution
$(y,Y)\in L^{2}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R}))\times L^{2}_{\mathbb{F}}(0,T;\mathbb{R}).$ (113)
We will show that one can choose a particular terminal condition $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;\mathbb{R})$ such that the corresponding solution satisfies $y\in L^{\infty}_{\mathbb{F}}(0,T;\mathbb{R})$ while ensuring that $Y\notin L^{\infty}_{\mathcal{F}}(\Omega;L^{2}(0,T;\mathbb{R}))$.
Consider $y_{T}=\mathbf{1}_{\{W_{T}\geq c\}}$ where $c>0$ is a fixed number and where we use the notation $W_{\cdot}:=W(\cdot)$. Note that by construction $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;\mathbb{R})$. Moreover, by using the change of variables $(z,Z):=e^{-\lambda t}(y,Y)$ and Itô’s formula, equation (112) becomes
$\begin{cases}\textnormal{d}z=Z\,\textnormal{d}W(t)\quad t\in(0,T),\\ z(T)=e^{-\lambda T}\mathbf{1}_{\{W_{T}\geq c\}}.\end{cases}$ (114)
By the martingale representation theorem, it is not difficult to see that the process
$z(t)={\mathbb{E}}\left(e^{-\lambda T}\mathbf{1}_{\{W_{T}\geq c\}}\;{\Big|}\;\mathcal{F}_{t}\right),\quad t\in(0,T),$ (115)
solves equation (114) (together with its initial condition). Moreover, by the change of variable $y=e^{\lambda t}z$, $y(T)=\mathbf{1}_{\{W_{T}\geq c\}}$ and by independence of the Brownian increments
$y(t)$ $=e^{\lambda t}{\mathbb{E}}\left(e^{-\lambda T}\mathbf{1}_{\{W_{T}\geq c\}}\;{\Big|}\;\mathcal{F}_{t}\right)=e^{-\lambda(T-t)}\mathbb{P}\left(W_{T-t}\geq c-W_{t}\right)$
$=e^{-\lambda(T-t)}\int_{\mathbb{R}}\mathbb{P}\left(W_{T-t}\geq c-x\;{\Big|}\;W_{t}=x\right)\mathbb{P}(W_{t}\in\textnormal{d}x),$
whence,
$y(t)=e^{-\lambda(T-t)}\Phi\left(\frac{c-W_{t}}{\sqrt{T-t}}\right),$ (116)
where $\Phi(x):=\displaystyle\int_{x}^{\infty}\frac{e^{-u^{2}/2}}{\sqrt{2\pi}}\textnormal{d}{u}$. Since $\|\Phi\|_{L^{\infty}(\mathbb{R})}=\frac{1}{\sqrt{2\pi}}$ it follows that $y\in L^{\infty}(\Omega;L^{\infty}(0,T;\mathbb{R}))$. Moreover, $z$ is adapted to $\mathbb{F}$ by construction (see (115)), hence $y=e^{\lambda t}z$ is adapted as well, and therefore $y\in L^{\infty}_{\mathbb{F}}(0,T;\mathbb{R})$.
Let us define $v(t):=\frac{c-W_{t}}{\sqrt{T-t}}$. Applying Itô’s formula to (116) yields after a straightforward computation that
$\textnormal{d}y(t)=\lambda e^{-\lambda(T-t)}\Phi(v(t))\textnormal{d}t-\frac{e^{-\lambda(T-t)}}{\sqrt{T-t}}\Phi^{\prime}(v(t))\textnormal{d}{W(t)}.$ (117)
Comparing this with (112) and (116), we identify for $t\in(0,T)$
$Y(t)=\frac{1}{\sqrt{2\pi}}\frac{e^{-\lambda(T-t)}}{\sqrt{T-t}}\exp\left(-\frac{1}{2}\left(\frac{c-W_{t}}{\sqrt{T-t}}\right)^{2}\right),\quad\text{a.s.}$ (118)
Proposition 6.1.
The process $Y$ defined (118) does not belong to $L^{\infty}_{\mathbb{F}}(\Omega;L^{2}(0,T;\mathbb{R}))$.
Proof.
Let us define the function
$\omega\in\Omega\mapsto I(\omega):=\int_{0}^{T}\frac{1}{2\pi}\frac{e^{-2\lambda(T-t)}}{{T-t}}\exp\left(-\left(\frac{c-W_{t}(\omega)}{\sqrt{T-t}}\right)^{2}\right)\textnormal{d}t.$
Note that $I(\omega)\geq 0$ and from (113), $I(\omega)<+\infty$ for almost all $\omega\in\Omega$. So, to prove that $Y\notin L^{\infty}(\Omega;L^{2}(0,T))$, it is enough to show that for every given constant $M>0$
$\mathbb{P}\left(I\geq M\right)>0.$
Let $\delta\in(0,T/4)$ be a given and consider the intervals
$I_{j}=\left[T-2^{-j}\delta,T-2^{-j-1}\delta\right]$
so that $\cup_{j\in\mathbb{N}}I_{j}=[T-\delta,T]$. We note that $|I_{j}|=\delta 2^{-j-1}$. Let $\epsilon>0$ be a fixed quantity and define the events
$A_{j}=\left\{\sup_{t\in I_{j}}\left|W_{t}-W_{T-\delta 2^{-j}}\right|\leq\epsilon\sqrt{|I_{j}|}\quad\text{and}\quad|W_{T-\delta 2^{-j}}-c|\leq\epsilon\sqrt{|I_{j}|}\right\}.$
We have the following intermediate result.
Lemma 6.2.
Let $j_{0}\in\mathbb{N}$ be large enough. Define $A^{N}:=\cap_{j=j_{0}}^{N}A_{j}$ for some natural number $N>j_{0}$. Then
$\mathbb{P}(A^{N})>0.$ (119)
We postpone the proof of this result to the end of this section. Assuming that Lemma 6.2 holds, we have that
$I(\omega)$ $\geq\frac{e^{-2\lambda T}}{2\pi}\int_{\cup_{j=j_{0}}^{N}I_{j}}\frac{1}{{T-t}}\exp\left(-\left(\frac{c-W_{t}(\omega)}{\sqrt{T-t}}\right)^{2}\right)\textnormal{d}t.$ (120)
Moreover, if $\omega\in A^{N}$, then for all $t\in I_{j}$ with $j\in\{j_{0},\ldots,N\}$
$|W_{t}(\omega)-c|\leq\sup_{t\in I_{j}}\left|W_{t}(\omega)-W_{T-\delta 2^{-j}}(\omega)\right|+|W_{T-\delta 2^{-j}}(\omega)-c|\leq 2\epsilon\sqrt{|I_{j}|},$ (121)
and
$(T-t)^{-1}\leq 2^{j+1}\delta^{-1}.$ (122)
Putting together estimates (120) to (122), we get
$I(\omega)\geq\frac{e^{-2\lambda T}}{2\pi}e^{-4\epsilon^{2}}\sum_{j=j_{0}}^{N}\int_{I_{j}}\frac{1}{T-t}\textnormal{d}t=\frac{e^{-2\lambda T}}{2\pi}e^{-4\epsilon^{2}}(N-j_{0})\ln(2).$
So, given any constant $M>0$, we can always find a finite number $N>0$ such that
$I(\omega)\geq M\quad\text{for all $\omega\in A^{N}$}.$
Therefore, $\mathbb{P}(I\geq M)>0$ from (119) and this concludes the proof. ∎
Remark 6.3.
Note that the probability of the events $A^{N}$ appearing in (119) may become very small as $N$ increases. However, since $N$ is finite, this probability is always strictly positive. This is consistent with the fact that $I<+\infty$ a.s. and highlights that one cannot obtain any uniform bound for the trajectories in $\Omega$.
We conclude this section with the following.
Proof of Lemma 6.2.
For simplicity, for all $i\in\{j_{0},\ldots,N\}$, we set
$B_{i}$ $:=\left\{\sup_{t\in I_{i}}\left|W_{t}-W_{T-\delta 2^{-i}}\right|\leq\varepsilon\sqrt{|I_{i}|}\right\}$ (123)
therefore $A_{i}:=\left\{\left|W_{T-\delta 2^{-i}}-c\right|\leq\varepsilon\sqrt{|I_{i}|}\right\}\cap B_{i}.$
Let $t_{i}:=T-\delta 2^{-i}$ for $i\in\{j_{0},\ldots,N\}$, and define the function
$\psi_{i}(x):=\mathbb{P}\!\left(\bigcap_{k=i}^{N}A_{k}\,\middle|\,W_{t_{i}}=x\right),\qquad i=j_{0},\ldots,N.$ (124)
The conditioning is understood in the sense of regular conditional probabilities, which are well defined for almost every $x\in\mathbb{R}$ with respect to the law of $W_{t_{i}}$. The function $\psi_{i}$ represents the probability that all remaining events from $i$ to $N$ are satisfied, given the position of the Brownian motion at time $t_{i}$. This interpretation allows us to derive the following recursive representation
$\psi_{i}(x)$ $=\mathbb{E}\!\left(\mathbf{1}_{\bigl\{|W_{t_{i}}-c|\leq\varepsilon\sqrt{|I_{i}|}\bigr\}}\,\mathbf{1}_{B_{i}}\,\mathbf{1}_{\bigcap_{k=i+1}^{N}A_{k}}\;\middle|\;W_{t_{i}}=x\right)$
$=\mathbf{1}_{\bigl\{|x-c|\leq\varepsilon\sqrt{|I_{i}|}\bigr\}}\,\mathbb{E}\!\left(\mathbf{1}_{B_{i}}\,\mathbf{1}_{\bigcap_{k=i+1}^{N}A_{k}}\;\middle|\;W_{t_{i}}=x\right).$ (125)
The identity (6) follows by noting that, under the conditioning $W_{t_{i}}=x$, the constraint on $W_{t_{i}}$ becomes deterministic. Finally, exploiting the tower property of conditional expectation and the Markov property, we get
$\psi_{i}(x)$ $=\mathbf{1}_{\bigl\{|x-c|\leq\varepsilon\sqrt{|I_{i}|}\bigr\}}\mathbb{E}\!\left(\mathbf{1}_{B_{i}}\,\psi_{i+1}(W_{t_{i+1}})\,\middle|\,W_{t_{i}}=x\right),\qquad i=j_{0},\ldots,N-1.$
To obtain the desired result, we look for a strictly positive lower bound for $\psi_{i}$. Observe that if $|x-c|>\varepsilon\sqrt{|I_{i}|}$, then $\psi_{i}(x)=0$. For this reason, it suffices to restrict our attention to $x$ close to $c$ and introduce the interval
$\tilde{I}_{i}:=\left[\,c-\frac{\varepsilon\sqrt{|I_{i}|}}{4},\;c+\frac{\varepsilon\sqrt{|I_{i}|}}{4}\,\right].$ (126)
Observe that if $x\in\tilde{I}_{i}$, then $\mathbf{1}_{\bigl\{|x-c|\leq\varepsilon\sqrt{|I_{i}|}\bigr\}}=1$. We therefore define
$q_{i}:=\inf_{x\in\tilde{I}_{i}}\psi_{i}(x).$ (127)
We now proceed by backward induction to show that the quantities $q_{i}$ are strictly positive for all $i\in\{j_{0},\ldots,N\}$. For the terminal index $i=N$ and for all $x\in\tilde{I}_{N}$, we have $|x-c|\leq\varepsilon\sqrt{|I_{N}|}$. Hence, we note that
$\psi_{N}(x)=\mathbb{P}(B_{N}\mid W_{t_{N}}=x).$
Using the scaling property of Brownian motion we obtain
$\psi_{N}(x)$ $=\mathbb{P}\!\left(\sup_{t\in I_{N}}\left|W_{t}-W_{t_{N}}\right|\leq\varepsilon\sqrt{|I_{N}|}\right)$
$=\mathbb{P}\!\left(\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\right)=:\kappa_{\varepsilon},$ (128)
where
$\left\{\frac{W_{t_{N}+s|I_{N}|}-W_{t_{N}}}{\sqrt{|I_{N}|}}\right\}_{s\in[0,1]}\overset{d}{\sim}\{\widetilde{W}_{s}\}_{s\in[0,1]},$
and $\widetilde{W}$ is a standard Brownian motion. Here $\kappa_{\varepsilon}>0$ is a positive constant independent of $N$. This constant is simply the probability that a standard Brownian motion remains inside $\varepsilon,\varepsilon]$ throughout the time interval $[0,1]$. This is a classical fact concerning the distribution of the maximum of Brownian motion, see, for instance, [RES92, Section 6.7]. In particular, $q_{N}>0$.
Assume now that $q_{i+1}>0$ and let $x\in\tilde{I}_{i}$. We show that $q_{i}>0$. In more detail,
$\psi_{i}(x)$ $=\mathbb{E}\!\left(\mathbf{1}_{B_{i}}\,\psi_{i+1}(W_{t_{i+1}})\,\middle|\,W_{t_{i}}=x\right)$
$\geq q_{i+1}\,\mathbb{P}\!\left(B_{i}\cap\Big\{|W_{t_{i+1}}-c|\leq\tfrac{\varepsilon}{4\sqrt{2}}\sqrt{|I_{i}|}\Big\}\,\middle|\,W_{t_{i}}=x\right).$ (129)
Here we used the definition of $q_{i+1}$ together with the fact that $|I_{i+1}|=\tfrac{1}{2}|I_{i}|$. Indeed,
$|W_{t_{i+1}}-c|\leq\tfrac{\varepsilon}{4\sqrt{2}}\sqrt{|I_{i}|}=\tfrac{\varepsilon}{4}\sqrt{|I_{i+1}|}$
implies that $W_{t_{i+1}}\in\tilde{I}_{i+1}$, and therefore
$\psi_{i+1}(W_{t_{i+1}})\geq q_{i+1}.$
To bound the probability appearing in (129), we apply a rescaling argument analogous to that used in (128). Conditioning on $W_{t_{i}}=x$ and using the scaling property of Brownian motion, we obtain
$\mathbb{P}\!\left(B_{i}\cap\Big\{|W_{t_{i+1}}-c|\leq\tfrac{\varepsilon}{4\sqrt{2}}\sqrt{|I_{i}|}\Big\}\,\middle|\,W_{t_{i}}=x\right)$
$=\mathbb{P}\!\left(\left\{\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\right\}\cap\left\{|\widetilde{W}_{1}-m|\leq\tfrac{\varepsilon}{4\sqrt{2}}\right\}\right),$ (130)
where $m=\frac{c-x}{\sqrt{|I_{i}|}}$. Since $x\in\tilde{I}_{i}$, we have
$|m|=\frac{|c-x|}{\sqrt{|I_{i}|}}\leq\frac{\varepsilon}{4}.$
Consequently,
$\Big[-\tfrac{\varepsilon}{4\sqrt{2}}+m,\;\tfrac{\varepsilon}{4\sqrt{2}}+m\Big]\subset(-\varepsilon,\varepsilon).$
Therefore,
$\mathbb{P}\!\left(\left\{\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\right\}\cap\left\{|\widetilde{W}_{1}-m|\leq\tfrac{\varepsilon}{4\sqrt{2}}\right\}\right)$
$=\mathbb{P}\big(\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\big)\,\mathbb{P}\!\left(\widetilde{W}_{1}\in\Big[-\tfrac{\varepsilon}{4\sqrt{2}}+m,\,\tfrac{\varepsilon}{4\sqrt{2}}+m\Big]\,\middle|\,\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\right).$ (131)
Since the conditional law of $\widetilde{W}_{1}$ given $\{\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\}$ admits a continuous and strictly positive density on $(-\varepsilon,\varepsilon)$, and since all admissible intervals above are contained in a fixed compact subset of $(-\varepsilon,\varepsilon)$, there exists a constant $c_{\varepsilon}>0$, depending only on $\varepsilon$, such that
$\mathbb{P}\!\left(\widetilde{W}_{1}\in\Big[-\tfrac{\varepsilon}{4\sqrt{2}}+m,\,\tfrac{\varepsilon}{4\sqrt{2}}+m\Big]\,\middle|\,\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\right)\geq c_{\varepsilon},$
uniformly in $m$. Hence,
$\mathbb{P}\left(\left\{\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\right\}\cap\left\{|\widetilde{W}_{1}-m|\leq\tfrac{\varepsilon}{4\sqrt{2}}\right\}\right)\geq\kappa_{\varepsilon}\,c_{\varepsilon}>0.$
By the induction hypothesis $q_{i+1}>0$, so it follows that
$\psi_{i}(x)\geq q_{i+1}\kappa_{\varepsilon}\,c_{\varepsilon}>0\qquad\text{for all }x\in\tilde{I}_{i}.$
Since $q_{i}:=\inf_{x\in\tilde{I}_{i}}\psi_{i}(x)$, this proves that $q_{i}>0$. Therefore, $q_{i}>0$ for all $i\in\{j_{0},\ldots,N\}$, and
$q_{i}\geq q_{N}(\kappa_{\varepsilon}c_{\varepsilon})^{N-i}.$
Finally, by the law of total probability and the definition of the regular conditional probability $\psi_{j_{0}}$, we obtain
$\mathbb{P}\!\left(\bigcap_{k=j_{0}}^{N}A_{k}\right)$ $=\mathbb{E}\!\left(\psi_{j_{0}}(W_{t_{j_{0}}})\right)$
$\geq\mathbb{E}\!\left(\psi_{j_{0}}(W_{t_{j_{0}}})\mathbf{1}_{\{W_{t_{j_{0}}}\in\tilde{I}_{j_{0}}\}}\right)$
$\geq q_{N}(\kappa_{\varepsilon}c_{\varepsilon})^{N-j_{0}}\mathbb{P}\!\left(|W_{t_{j_{0}}}-c|\leq\tfrac{\varepsilon\sqrt{\delta\,2^{-j_{0}-1}}}{4}\right)$
$\geq q_{N}(\kappa_{\varepsilon}c_{\varepsilon})^{N-j_{0}}\tilde{\kappa}_{c,\varepsilon,T,\delta}\,\sqrt{|I_{j_{0}}|}>0,$
where $\tilde{\kappa}_{c,\varepsilon,T,\delta}>0$ is a constant depending only on the parameters $c$, $\varepsilon$, $T$, and $\delta$. This concludes the proof. ∎
Appendix A Auxiliary results
Lemma A.1 (Backward Grönwall inequality).
Let $g(t)$, $\alpha(t)$, $\beta(t)$ and $\gamma(t)$ be integrable functions with $\beta(t),\gamma(t)\geq 0$. For any $t\in[0,T]$, if
$g(t)\leq\alpha(t)+\beta(t)\int_{t}^{T}\gamma(s)g(s)\textnormal{d}{s},$
then
$g(t)\leq\alpha(t)+\beta(t)\int_{t}^{T}\alpha(u)\gamma(u)e^{\int_{t}^{u}\beta(s)\gamma(s)\textnormal{d}{s}}\textnormal{d}{u}.$
In particular, if $\alpha(t)\equiv\alpha$, $\beta(t)\equiv\beta$ and $\gamma(t)\equiv 1$, then
$g(t)\leq\alpha e^{\beta(T-t)}.$
Lemma A.2.
Let $(X,\mathcal{M},\mu)$ be a finite measure space and $f:X\to\mathbb{R}$ be a measurable function. We assume that there exists $K>0$ such that
$\forall p\geq 1,\ \left\lVert f\right\rVert_{p}\leq K.$ (132)
Then $f\in L^{\infty}(X)$ and
$\left\lVert f\right\rVert_{\infty}\leq K.$
Proof.
First, we prove that $f\in L^{\infty}(X)$ by proceeding as follows. We argue by contradiction. We assume that $\left\lVert f\right\rVert_{\infty}=+\infty$. Let $M>0$, and let us define $A_{M}=\{|f|\geq M\}$ then we have that $\mu(A_{M})>0$, then take $p$ large enough such that $\mu(A_{M})^{1/p}\geq 1/2$, then $\left\lVert f\right\rVert_{p}\geq(\mu(A_{M})M^{p})^{1/p}\geq M/2$ and since $M$ is arbitrary, this is contradiction. Thus, this proves that $f\in L^{\infty}(X)$.
Then, we prove the bound. This is a well-known fact that because $f\in L^{\infty}(X)$,
$\lim_{p\to+\infty}\left\lVert f\right\rVert_{p}=\left\lVert f\right\rVert_{\infty}.$
Therefore, by using the uniform bound on $L^{p}$ spaces of $f$ i.e. (132), we deduce the result. ∎
Lemma A.3.
Let $f\in C^{\infty}(\mathbb{R})$ and define $G(s)=\int_{0}^{1}(1-\sigma)f^{\prime\prime}(\sigma s)\textnormal{d}{\sigma}$. For any bounded closed interval $\mathcal{I}\subset\mathbb{R}$ containing the origin, there are constants $M,M_{1}>0$ only depending on $f$ and $\mathcal{I}$ such that
  1. i)
    $|G(s)|\leq M$ for all $s\in\mathcal{I}$.
  2. ii)
    For any $s_{1},s_{2}\in\mathcal{I}$, $|s_{1}^{2}G(s_{1})-s^{2}_{2}G(s_{2})|\leq M|s_{1}-s_{2}|\left(|s_{1}|+|s_{2}|\right)+M_{1}\left|s_{2}\right|^{2}|s_{1}-s_{2}|$.
Proof.
Point i) is direct with $M:=\max_{\tau\in\mathcal{I}}|f^{\prime\prime}(\tau)|$ since $f\in C^{\infty}(\mathbb{R})$ and $0\in\mathcal{I}$. For ii), let $s_{1},s_{2}\in\mathcal{I}$, then we can write
$s_{1}^{2}G(s_{1})-s^{2}_{2}G(s_{2})=\left(s_{1}^{2}-s_{2}^{2}\right)G(s_{1})+s_{2}^{2}\left(G(s_{1})-G(s_{2})\right)=:J_{1}+J_{2}.$ (133)
Using point i), we can estimate $J_{1}$ as follows
$|J_{1}|=\left|\left(s_{1}^{2}-s_{2}^{2}\right)\right||G(s_{1})|=\left|(s_{1}-s_{2})(s_{1}+s_{2})\right||G(s_{1})|\leq M|s_{1}-s_{2}|\left(|s_{1}|+|s_{2}|\right).$ (134)
On the other hand, w.l.o.g assume that $s_{1}>s_{2}$. For any $\sigma\in[0,1]$, note that since $f\in C^{\infty}(\mathbb{R})$, we have from mean value theorem that
$|f^{\prime\prime}(\sigma s_{1})-f^{\prime\prime}(\sigma s_{2})|=\left|\int_{\sigma s_{2}}^{\sigma s_{1}}f^{\prime}(u)\textnormal{d}{u}\right|\leq\sigma\max_{\tau\in\mathcal{I}}|f^{\prime}(\tau)||s_{1}-s_{2}|\leq M_{1}|s_{1}-s_{2}|,$ (135)
where we have used that $\sigma s_{2},\sigma s_{1}]\subset\mathcal{I}$ since $0\in\mathcal{I}$ and where $M_{1}:=\max_{\tau\in\mathcal{I}}|f^{\prime}(\tau)|$. Thus, from (135), we can estimate $J_{2}$ as
$|J_{2}|$ $=\left|s_{2}|^{2}\right|G(s-1)-G(s_{2})|=\left|s_{2}\right|^{2}\left|\int_{0}^{1}(1-\sigma)\left[f^{\prime\prime}(\sigma s_{1})-f^{\prime\prime}(\sigma s_{2})\right]\textnormal{d}{\sigma}\right|$
$\leq\left|s_{2}\right|^{2}\int_{0}^{1}|1-\sigma||f^{\prime\prime}(\sigma s_{1})-f^{\prime\prime}(\sigma s_{2})|\textnormal{d}{\sigma}\leq M_{1}\left|s_{2}\right|^{2}|s_{1}-s_{2}|.$ (136)
Putting together (133), (134) and (136) yields the desired result. ∎
Acknowledgements
The authors are grateful to the anonymous referees for their careful reading of the manuscript and for their valuable comments and suggestions, which helped improve the quality and clarity of the paper.
The authors would like to thank to LaSol (Solomon Lefschetz International Research Laboratory) for the financial support. Part of this work was done when the second author was visiting Instituto de Matemáticas, UNAM which provided excellent working conditions. The second author would like to thank Ying Hu for interesting discussions about this work.